### v6

# Survey of Hallucination in Natural Language Generation

Abstract.
Natural Language Generation (NLG) has improved exponentially in recent years thanks to the development of sequence-to-sequence deep learning technologies such as Transformer-based language models. This advancement has led to more fluent and coherent NLG, leading to improved development in downstream tasks such as abstractive summarization, dialogue generation and data-to-text generation. However, it is also apparent that deep learning based generation is prone to hallucinate unintended text, which degrades the system performance and fails to meet user expectations in many real-world scenarios. To address this issue, many studies have been presented in measuring and mitigating hallucinated texts, but these have never been reviewed in a comprehensive manner before.
In this survey, we thus provide a broad overview of the research progress and challenges in the hallucination problem in NLG. The survey is organized into two parts: (1) a general overview of metrics, mitigation methods, and future directions; (2) an overview of task-specific research progress on hallucinations in the following downstream tasks, namely abstractive summarization, dialogue generation, generative question answering, data-to-text generation, machine translation, and visual-language generation; and (3) hallucinations in large language models (LLMs)  [^1] This section was updated in Jan 2024.. This survey serves to facilitate collaborative efforts among researchers in tackling the challenge of hallucinated texts in NLG.
Hallucination, Intrinsic Hallucination, Extrinsic Hallucination, Faithfulness in NLG, Factuality in NLG, Consistency in NLG
copyright: nonejournalyear: 2022journal: CSURpublicationmonth: 2ccs: Computing methodologies Natural language generationccs: Computing methodologies Neural networks
1. Introduction
Natural Language Generation (NLG) is one of the crucial yet challenging sub-fields of Natural Language Processing (NLP). NLG techniques are used in many downstream tasks such as summarization, dialogue generation, generative question answering (GQA), data-to-text generation, and machine translation. Recently, the rapid development of NLG has captured the imagination of many thanks to the advances in deep learning technologies, especially Transformer (Vaswani et al., 2017)-based models like BERT (Devlin et al., 2019), BART (Lewis et al., 2020), GPT-2 (Radford et al., 2019), and GPT-3 (Brown et al., 2020). The conspicuous development of NLG tasks attracted the attention of many researchers, leading to an increased effort in the field.
Alongside the advancement of NLG models, attention towards their limitations and potential risks has also increased. Some early works (Welleck et al., 2019a; Holtzman et al., 2019) focus on the potential pitfalls of utilizing the standard likelihood maximization-based objective in training and decoding of NLG models. They discovered that such likelihood maximization approaches could result in degeneration, which refers generated output that is bland, incoherent, or gets stuck in repetitive loops. Concurrently, it is discovered that NLG models often generate text that is nonsensical, or unfaithful to the provided source input (Vinyals and Le, 2015; Koehn and Knowles, 2017a; Rohrbach et al., 2018; Raunak et al., 2021). Researchers started referring to such undesirable generation as hallucination (Maynez et al., 2020)  [^2] The term “hallucination” first appeared in Computer Vision (CV) in Baker and Kanade (2000) and carried more positive meanings, such as superresolution (Liu et al., 2007; Baker and Kanade, 2000), image inpainting (Fawzi et al., 2016), and image synthesizing (Zhang et al., 2019b). Such hallucination is something we take advantage of rather than avoid in CV. Nevertheless, recent works have started to refer to a specific type of error as ”hallucination” in image captioning (Biten et al., 2022; Rohrbach et al., 2018) and object detection (Kayhan et al., 2021; Bai et al., 2009), which denotes non-existing objects detected or localized incorrectly at their expected position. The latter conception is similar to “hallucination” in NLG..
Hallucination in NLG is concerning because it hinders performance and raises safety concerns for real-world applications. For instance, in medical applications, a hallucinatory summary generated from a patient information form could pose a risk to the patient. It may provoke a life-threatening incident for a patient if the instructions of a medicine generated by machine translation are hallucinatory. Hallucination can also lead to potential privacy violations. Carlini et al. (2020) demonstrate that language models can be prompted to recover and generate sensitive personal information from the training corpus (e.g., email address, phone/fax number, and physical address). Such memorization and recovery of the training corpus is considered a form of hallucination because the model is generating text that is not “faithful” to the source input content (i.e., such private information does not exist in the source input).
Currently, there are many active efforts to address hallucination for various NLG tasks. Analyzing hallucinatory content in different NLG tasks and investigating their relationship would strengthen our understanding of this phenomenon and encourage the unification of efforts from different NLG fields. However, to date, little has been done to understand hallucinations from a broader perspective that encompasses all major NLG tasks. To the best of our knowledge, existing surveys have only focused specific tasks like abstractive summarization (Maynez et al., 2020; Huang et al., 2021) and translation (Lee et al., 2019). Thus, in this paper, we present a survey of the research progress and challenges in the hallucination problem in NLG. And offer a comprehensive analysis of existing research on the phenomenon of hallucination in different NLG tasks, namely abstractive summarization, dialogue generation, generative question answering, data-to-text generation, and machine translation. We mainly discussed hallucination of the unimodal NLG tasks that have textual input sources upon which the generated text can be assessed. We also briefly summarize hallucinations in multi-modal settings such as visual-language tasks (Biten et al., 2022; Alayrac et al., 2022). This survey can provide researchers with a high-level insight derived from the similarities and differences of different approaches. Furthermore, given the various stages of development in studying hallucination from different tasks, the survey can assist researchers in drawing inspiration on concepts, metrics, and mitigation methods.
Organization of this Survey
The remainder of this survey is organized as follows. Section 2 $\sim$ Section 6 provide an overview of the hallucination problem in NLG by discussing the definition and categorization, contributors, metrics, and mitigation methods of hallucinations, respectively. The second part of our survey discusses the hallucination problem associated with specific NLG tasks: abstractive summarization in Section 7, dialogue generation in Section 8, GQA in Section 9, data-to-text generation in Section 10, machine translation in Section 11, and VL generation in Section 12. The third part discusses this phenomenon in LLMs in Section 13. Finally, we conclude the whole survey in Section 14.
2. Definitions
In the general context outside of NLP, hallucination is a psychological term referring to a particular type of perception (Fish et al., 2009; Macpherson and Platchias, 2013). Blom ([n. d.]) define hallucination as “a percept, experienced by a waking individual, in the absence of an appropriate stimulus from the extracorporeal world”. Simply put, a hallucination is an unreal perception that feels real. The undesired phenomenon of “NLG models generating unfaithful or nonsensical text” shares similar characteristics with such psychological hallucinations – explaining the choice of terminology. Hallucinated text gives the impression of being fluent and natural despite being unfaithful and nonsensical. It appears to be grounded in the real context provided, although it is actually hard to specify or verify the existence of such contexts. Similar to psychological hallucination, which is hard to tell apart from other “real” perceptions, hallucinated text is also hard to capture at first glance.
Within the context of NLP, the above definition of hallucination, the generated content that is nonsensical or unfaithful to the provided source content (Filippova, 2020; Maynez et al., 2020; Parikh et al., 2020; Zhou et al., 2021b), is the most inclusive and standard. However, there do exist variations in definition across NLG tasks, which will be further described in the later task-specific sections.
2.1. Categorization
Following the categorization from previous works (Maynez et al., 2020; Huang et al., 2021; Dziri et al., 2021a), there are two main types of hallucinations, namely intrinsic hallucination and extrinsic hallucination. To explain the definition and categorization more intuitively, we give examples of each category of hallucinations for each NLG downstream task in Table 1.
  1. (1)
    Intrinsic Hallucinations: The generated output that contradicts the source content. For instance, in the abstractive summarization task from Table 1, the generated summary “The first Ebola vaccine was approved in 2021” contradicts the source content “The first vaccine for Ebola was approved by the FDA in 2019.”.
  2. (2)
    Extrinsic Hallucinations: The generated output that cannot be verified from the source content (i.e., output that can neither be supported nor contradicted by the source). For example, in the abstractive summarization task from Table 1, the information “China has already started clinical trials of the COVID-19 vaccine.” is not mentioned in source. We can neither find evidence for the generated output from the source nor assert that it is wrong. Notably, the extrinsic hallucination is not always erroneous because it could be from factually correct external information (Maynez et al., 2020; Thomson and Reiter, 2020). Such factual hallucination can be helpful because it recalls additional background knowledge to improve the informativeness of the generated text. However, in most of the literature, extrinsic hallucination is still treated with caution because its unverifiable aspect of this additional information increases the risk from a factual safety perspective.
2.2. Task Comparison
The previous subsection is about the definition and categorization of hallucination commonly shared by many NLG tasks. Yet, there are some task-specific differences.
For the abstractive summarization, data-to-text, and dialogue tasks, the main difference is in what serves as the “source” and the level of tolerance towards hallucinations. The source in abstractive summarization is the input source text that is being summarized (See et al., 2017), while the source in data-to-text is non-linguistic data (Reiter and Dale, 1997; Gatt and Krahmer, 2018), and the source(s) in the dialogue system is dialogue history and/or the external knowledge sentences. Tolerance towards hallucinations is very low in both the summarization (Pagnoni et al., 2021) and data-to-text tasks (Parikh et al., 2020; Wang et al., 2021a, 2020b) because it is essential to provide faithful generation. In contrast, the tolerance is relatively higher in dialogue systems because the desired characteristics are not only faithfulness but also user engagement, especially in open-domain dialogue systems (Huang et al., 2020b; Ji et al., 2022).
For the generative question answering (GQA) task, the exploration of hallucination is at its early stage, so there is no standard definition or categorization of hallucination yet. However, we can see that the GQA literature mainly focuses on “intrinsic hallucination” where the source is the world knowledge (Li et al., 2021a). Lastly, unlike the aforementioned tasks, the categorizations of hallucinations in machine translation vary within the task. Most relevant literature agrees that translated text is considered a hallucination when the source text is completely disconnected from the translated target (Lee et al., 2019; Müller et al., 2020; Raunak et al., 2021). For further details, please refer to Section 11.
2.3. Terminology Clarification
Multiple terminologies are associated with the concept of hallucination. We provide clarification of the commonly used terminologies hallucination, faithfulness, and factuality to resolve any confusion. Faithfulness is defined as staying consistent and truthful to the provided source – an antonym to ”hallucination.” Any work that tries to maximize faithfulness thus focuses on minimizing hallucination. For this reason, our survey includes all those works that address the faithfulness of machine-generated outputs. Factuality refers to the quality of being actual or based on fact. Depending on what serves as the “fact”, ”factuality” and ”faithfulness” may or may not be the same. Maynez et al. (2020) differentiate ”factuality” from ”faithfulness” by defining the “fact” to be the world knowledge. In contrast, Dong et al. (2020) use the source input as the “fact” to determine the factual correctness, making ”factuality” indistinguishable from ”faithfulness”. In this paper, we adopt the definition from Maynez et al. (2020) because we believe having such a distinction between source knowledge and world knowledge provides a more clear understanding.
Note that the judging criteria for what is considered faithful or hallucinated (i.e., the definition of hallucination) can differ across tasks. For more details of these variation definitions, you can find in the later task-specific sections.
Task Sub-Task Type Source Output
Abstractive Summarization  (Pagnoni et al., 2021) Intrinsic
The first vaccine for Ebola was approved by the FDA in 2019 in the US, five years after the initial outbreak in 2014. To produce the vaccine, scientists had to sequence the DNA
The first Ebola vaccine was approved in 2021.
Extrinsic
of Ebola, then identify possible vaccines, and finally show successful clinical trials. Scientists say a vaccine for COVID-19 is unlikely to be ready this year, although clinical trials have already started.
China has already started clinical trials of the COVID-19 vaccine.
Dialogue Task- oriented  (Li et al., 2020c) Intrinsic inform (NAME = pickwick hotel, PRICERANGE = moderate)
the hotel named pickwick hotel is in a high price range.
Extrinsic
the pickwick hotel in san diego is a moderate price range
Open- domain (Santhanam et al., 2021) Intrinsic
Dialog History:
Speaker 1: What do you think about Murray?
Speaker 2: I think Murray is a great player he just needs to stay healthy in order to compete more. Who do you like best?
Speaker 1: I like Roger Nadal the most!
Extrinsic
Knowledge: In tennis, the term Big Four refers to the quartet of men’s singles players comprising Roger Federer, Rafael Nadal, Novak Djokovic, and Andy Murray.
Speaker 1: I like Djokovic. He has played in the top ten singles players of the world.
Generative QA  (Nguyen et al., 2016) Intrinsic
Question: dow jones industrial average please?
Answer: The Dow Jones Industrial Average(DJIA) is an index of 30 major U.S.stock indexes. The DJIAwas created in 1896 by Charles Dow, a statistician and editor of The Wall Street Journal. The index was first published in the mid-1880s, and was last updated on May 26,1896.
Extrinsic Question: definition of sadduction
Answer: Sadducees. noun. The definition of a sadducee is a person who acts in a deceitful or duplicitous manner. An example of a Sadduceee is a politician who acts deceitfully in order to gain political power.1 a member of a Jewish sect that was active during the Second Temple.
Data2text  (Wiseman et al., 2017) Intrinsic TEAM CITY WIN LOSS PTS FG_PCT BLK The Houston Rockets (18-4) defeated the Denver Nuggets (10-13) 108-96 on Saturday.
Rockets Houston 18 5 108 44 7
Extrinsic Nuggets Denver 10 13 96 38 7
Houston has won two straight games and six of their last seven.
Translation  (Zhou et al., 2021b) Intrinsic
迈克周四去书店。(Michael went to the bookstore on Thursday.)
Jerry didn’t go to the bookstore.
Extrinsic
迈克周四去书店。(Michael went to the bookstore on Thursday.)
Michael happily went to the bookstore with his friend.
Table 1. Examples of each category of hallucinations for each task. In the Data2Text task, H/A: home/away, MIN: minutes, PTS: points, REB: rebounds, AST: assists, BLK: blocks, FG_PCT: field goals percentage. The examples for VL tasks are shown in Figure 2 and Figure 3
Category Task Works
Automatic Metrics Statistical Dialogue Shuster et al. (2021)
Data2Text Dhingra et al. (2019); Wang et al. (2020b)
Translation Martindale et al. (2019)
Captioning Rohrbach et al. (2018)
Model- based
Abstractive
Summarization
Nan et al. (2021); Durmus et al. (2020); Wang et al. (2020a); Kryscinski et al. (2020)
Goodrich et al. (2019); Pagnoni et al. (2021); Gabriel et al. (2021); Zhou et al. (2021b)
Laban et al. (2022); Falke et al. (2019); Mishra et al. (2021); Scialom et al. (2021)
Dialogue
Li et al. (2020c); Balakrishnan et al. (2019); Honovich et al. (2021)
Dziri et al. (2021b); Gupta et al. (2021); Santhanam et al. (2021)
Generative QA
Sellam et al. (2020),Zhang et al. (2020a),Durmus et al. (2020)
Wang et al. (2020a), Su et al. (2022)
Data2Text
Liu et al. (2021b); Wiseman et al. (2017); Dušek and Kasner (2020)
Tian et al. (2020); Filippova (2020); Rebuffel et al. (2021)
Translation
Kong et al. (2019); Lee et al. (2019); Tu et al. (2016)
Feng et al. (2020); Garg et al. (2019); Zhou et al. (2021b)
Raunak et al. (2021); Parthasarathi et al. (2021)
Task-Agnostic Goyal and Durrett (2020); Zhou et al. (2021b); Liu et al. (2021a)
Mitigation Method Data- Related
Abstractive
Summarization
Nan et al. (2021); Cao et al. (2018); Zhu et al. (2021)
Gunel et al. (2019)
Dialogue
Shen et al. (2021); Wu et al. (2021); Honovich et al. (2021)
Santhanam et al. (2021); Shuster et al. (2021)
Generative QA Yin et al. (2016); Bi et al. (2019); Fan et al. (2019a)
Data2Text
Parikh et al. (2020); Wang (2019); Nie et al. (2019); Liu et al. (2021b)
Rebuffel et al. (2022); Nie et al. (2018)
Translation
Raunak et al. (2021); Lee et al. (2019)
Junczys-Dowmunt (2018); Briakou and Carpuat (2021)
Captioning Biten et al. (2022)
Modeling and Inference
Abstractive
Summarization
Huang et al. (2020a); Li et al. (2018); Song et al. (2020a)
Cao et al. (2020); Aralikatte et al. (2021); Cao and Wang (2021)
Chen et al. (2021b); Albrecht and Hwa (2007); Zhao et al. (2020)
Dialogue
Rashkin et al. (2021); Balakrishnan et al. (2019); Li et al. (2020c)
Dziri et al. (2021a)
Generative QA
Li et al. (2021a); Krishna et al. (2021); Fan et al. (2019a)
Nakano et al. (2021); Su et al. (2022)
Data2Text
Liu et al. (2021b); Xu et al. (2021a); Tian et al. (2020); Wang et al. (2021a, 2020b)
Su et al. (2021); Rebuffel et al. (2022); Filippova (2020); Xiao and Wang (2021)
Puduppully and Lapata (2021)
Translation
Feng et al. (2020); Weng et al. (2020); Lee et al. (2019)
Li et al. (2020a); Raunak et al. (2021); Wang and Sennrich (2020)
Zhou et al. (2021b); Bengio et al. (2015)
Goyal et al. (2017); Xu et al. (2019)
Captioning Xiao and Wang (2021); Dai et al. (2022b)
Table 2. Evaluation metrics and mitigation methods for each task. *The hallucination metrics are not specifically proposed for generative question answering (GQA), but they can be adapted for that task.
3. Contributors to Hallucination in NLG
3.1. Hallucination from Data
The main cause of hallucination from data is source-reference divergence. This divergence happens 1) as an artifact of heuristic data collection or 2) due to the nature of some NLG tasks that inevitably contain such divergence. When a model is trained on data with source-reference(target) divergence, the model can be encouraged to generate text that is not necessarily grounded and not faithful to the provided source.
Heuristic data collection
When collecting large-scale datasets, some works heuristically select and pair real sentences or tables as the source and target (Lebret et al., 2016; Wiseman et al., 2017). As a result, the target reference may contain information that cannot be supported by the source (Wang, 2019; Parikh et al., 2020). For instance, when constructing WIKIBIO (Lebret et al., 2016), a dataset for generating biographical notes based on the infoboxes of Wikipedia, the authors took the Wikipedia infobox as the source and the first sentence of the Wikipedia page as the target ground-truth reference. However, the first sentence of the Wikipedia article is not necessarily equivalent to the infobox in terms of the information they contain. Indeed, Dhingra et al. (2019) points out that 62% of the first sentences in WIKIBIO have additional information not stated in the corresponding infobox. Such mismatch between source and target in datasets can lead to hallucination.
Another problematic scenario is when duplicates from the dataset are not properly filtered out. It is almost impossible to check hundreds of gigabytes of text corpora manually. Lee et al. (2021) show that duplicated examples from the pretraining corpus bias the model to favor generating repeats of the memorized phrases from the duplicated examples.
Innate divergence
Some NLG tasks by nature do not always have factual knowledge alignment between the source input text and the target reference, especially those that value diversity in generated output. For instance, it is acceptable for open-domain dialogue systems to respond in chit-chat style, subjective style (Rashkin et al., 2021), or with a relevant fact that is not necessarily present in the user input, history or provided knowledge source – this improves the engagingness and diversity of the dialogue generation. However, researchers have discovered that such dataset characteristic leads to inevitable extrinsic hallucinations.
3.2. Hallucination from Training and Inference
As discussed in the previous subsection, source-reference divergence existing in dataset is one of the contributors of hallucination. However, Parikh et al. (2020) show that hallucination problem still occurs even when there is very little divergence in dataset. This is because there is another contributor of hallucinations – training and modeling choices of neural models (Vinyals and Le, 2015; Koehn and Knowles, 2017a; Rohrbach et al., 2018; Raunak et al., 2021).
Imperfect representation learning
The encoder has the role of comprehending and encoding input text into meaningful representations. An encoder with a defective comprehension ability could influence the degree of hallucination (Parikh et al., 2020). When encoders learn wrong correlations between different parts of the training data, it could result in erroneous generation that diverges from the input (Tian et al., 2020; Feng et al., 2020; Aralikatte et al., 2021; Li et al., 2018).
Erroneous decoding
The decoder takes the encoded input from the encoder and generates the final target sequence. Two aspects of decoding contribute to hallucinations. First, decoders can attend to the wrong part of the encoded input source, leading to erroneous generation (Tian et al., 2020). Such wrong association results in generation with facts mixed up between two similar entities (Shuster et al., 2021; Dziri et al., 2021a). Second, the design of the decoding strategy itself can contribute to hallucinations. Dziri et al. (2021a) illustrate that a decoding strategy that improves the generation diversity, such as top-k sampling, is positively correlated with increased hallucination. We conjecture that deliberately added “randomness” by sampling from the top-k samples instead of choosing the most probable token increase the unexpected nature of the generation, leading to a higher chance of containing hallucinated content.
Exposure Bias
Regardless of decoding strategy choices, the exposure bias problem (Bengio et al., 2015; Ranzato et al., 2016), defined as the discrepancy in decoding between training and inference time, can be another contributor to hallucination. It is common practice to train the decoder with teacher-forced maximum likelihood estimation (MLE) training, where the decoder is encouraged to predict the next token conditioned on the ground-truth prefix sequences. However, during the inference generation, the model generates the next token conditioned on the historical sequences previously generated by itself (He et al., 2021). Such a discrepancy can lead to increasingly erroneous generation, especially when the target sequence gets longer.
Parametric knowledge bias
Pre-training of models on a large corpus is known to result in the model memorizing knowledge in its parameters (Petroni et al., 2019; Roberts et al., 2020; Madotto et al., 2020b). This so-called parametric knowledge helps improve the performance of downstream tasks but also serves as another contributor to hallucinatory generation. Large pre-trained models used for downstream NLG tasks are powerful in providing generalizability and coverage, but Longpre et al. (2021) have discovered that such models prioritize parametric knowledge over the provided input. In other words, models that favor generating output with their parametric knowledge instead of the information from the input source can result in the hallucination of excess information in the output. On the other hand, current research works (Yin et al., 2023a; Burns et al., 2022; Rajpurkar et al., 2018; Kadavath et al., 2022; Manakul et al., 2023) highlight a discrepancy between surface realization and inherent knowledge of the model in NLG tasks. Models can realize they are generating something hallucinated in some way.
4. Metrics Measuring Hallucination
Recently, various studies have illustrated that most conventional metrics used to measure the quality of writing are not adequate for quantifying the level of hallucination (Reiter, 2018). It has been shown that state-of-the-art abstractive summarization systems, evaluated with metrics such as ROUGE, BLEU, and METEOR, have hallucinated content in 25% of their generated summaries (Falke et al., 2019). A similar phenomenon has been shown in other NLG tasks, where it has been discovered that traditional metrics have a poor correlation with human judgment in terms of the hallucination problem (Krishna et al., 2021; Dhingra et al., 2019; Durmus et al., 2020; Honovich et al., 2021). Therefore, there are active research efforts to define effective metrics for quantifying hallucination. FRANK (Pagnoni et al., 2021) surveys the faithfulness metrics for summarization and compares these metrics’ correlations with human judgments. To assess the example-level accuracy of metrics in diverse tasks, TRUE (Honovich et al., 2022) reports their Area Under the ROC Curve (ROC AUC) in regard to hallucinated example detection.
4.1. Statistical Metric
One of the simplest approaches is to leverage lexical features (n-grams) to calculate the information overlap and contradictions between the generated and the reference texts – the higher the mismatch counts, the lower the faithfulness and thus the higher the hallucination score.
Given that many traditional metrics leverage the target text as the ground-truth reference (e.g., ROUGE, BLEU, etc.), Dhingra et al. (2019) build upon this idea and propose PARENT (Precision And Recall of Entailed n-grams from the Table)  [^3] Note that PARENT is a general metric like ROUGE and BLEU, not only constrained to hallucination, a metric which can also measure hallucinations using both the source and target text as references. Particularly, PARENT n-gram lexical entailment matches generated text with both the source table and target text. The F1-score that combines the precision and recall of the entailment reflects the accuracy in the table-to-text task. The source text is additionally used because it is not guaranteed that the output target text contains the complete set of information available in the input source text.
It is common for NLG tasks to have multiple plausible outputs from the same input, which is known as one-to-many mapping (Su et al., 2020; Guan and Huang, 2020). In practice, however, covering all the possible outputs is too expensive and almost impossible. Thus, many works simplify the hallucination evaluation setup by relying on the source text as the sole reference. Their metrics just focus on the information referred by input sources to measure hallucinations, especially intrinsic hallucinations. For instance, Wang et al. (2020b) propose PARENT-T, which simplifies PARENT by only using table content as the reference. Similarly, Knowledge F1 (Shuster et al., 2021) – a variant of unigram F1 – has been proposed for knowledge-grounded dialogue tasks to measure the overlap between the model’s generation and the knowledge used to ground the dialogue during dataset collection.
Furthermore, Martindale et al. (2019) proposed a bag-of-vectors sentence similarity (BVSS) metric for measuring sentence adequacy in machine translation, that only refers to the target text. This statistical metric helps to determine whether the MT output has a different amount of information than the translation reference.
Although simple and effective, one potential limitation of the lexical matching is that it can only handle the lexical information. Thus, it fails to deal with syntactic or semantic variations (Sellam et al., 2020).
4.2. Model-based Metric
Model-based metrics leverage neural models to measure the hallucination degree in the generated text. They are proposed to handle more complex syntactic and even semantic variations. The model-based metrics comprehend the source and generated texts and detect the knowledge/content mismatches. However, the neural models can be subject to errors that can propagate and adversely affect the accurate quantification of hallucination.
4.2.1. Information Extraction (IE)-based
It is not always easy to determine which part of the generated text contains the knowledge that requires verification. IE-based metrics use IE models to represent the knowledge in a simpler relational tuple format (e.g., subject, relation, object), then verify against relation tuples extracted from the source/reference. Here, the IE model is identifying and extracting the “facts” that require verification. In this way, words containing no verifiable information (e.g., stopwords, conjunctions, etc) are not included in the verification step.
[…]
4.2.2. QA-based
This approach implicitly measures the knowledge overlap or consistency between the generation and the source reference. This is based on the intuition that similar answers will be generated from a same question if the generation is factually consistent with the source reference. It is already put in use to evaluate hallucinations in many tasks, such as summarization  (Wang et al., 2020a; Durmus et al., 2020; Scialom et al., 2021), dialogue (Honovich et al., 2021), and data2text generation (Rebuffel et al., 2021).
QA-based metric that measures the faithfulness of the generated text is consisted of three parts: First, given a generated text, a question generation (QG) model generates a set of question-answer pairs. Second, a question answering (QA) model answers the generated questions given a ground-truth source text as the reference (containing knowledge). Lastly, the hallucination score is computed based on the similarity of the corresponding answers.
Similar to the IE-based metrics, the limitation of this approach is the potential error that might arise and propagated from either the QG model or the QA model.
4.2.3. Natural Language Inference (NLI) Metrics
There are not many labelled datasets for hallucination detection tasks, especially at the early stage when the hallucination problem starts to gain attention. As an alternative, many works leverage the NLI dataset to tackle hallucinations. Note that NLI is a task that determines whether a “hypothesis” is true (entailment), false (contradiction), or undetermined (neutral) given a “premise”. These metrics are based on the idea that only the source knowledge reference should entail the entirety of the information in faithful and hallucination-free generation  (Williams et al., 2018; Falke et al., 2019; Mishra et al., 2021; Kryscinski et al., 2020; Dušek and Kasner, 2020; Huang et al., 2021; Honovich et al., 2021; Laban et al., 2022; Dziri et al., 2021b). More specifically, NLI-based metrics define the hallucination/faithfulness score to be the entailment probability between the source and its generated text, also known as the percentage of times generated text entails, neutral to, and contradicts the source.
According to Honovich et al. (2021), NLI-based approaches are more robust to lexical variability than token matching approaches such as IE-based and QA-based metrics. Nevertheless, as illustrated by Falke et al. (2019), off-the-shelf NLI models tend to transfer poorly to the abstractive summarization task. Thus, there is a line of research in improving and extending the NLI paradigm specifically for hallucination evaluation purposes (Falke et al., 2019; Dziri et al., 2021b). Apart from generalizability, Goyal and Durrett (2020) point out the potential limitation of using sentence-level entailment models, namely their incapability to pinpoint and locate which parts of the generation are erroneous. In response, the authors propose a new dependency-level entailment and attempt to identify factual inconsistencies in a more fine-grained manner.
4.2.4. Faithfulness Classification Metrics
To improve upon NLI-based metrics, task-specific datasets are constructed to improve from the NLI-based metrics. Zhou et al. (2021b); Liu et al. (2021a) constructed syntactic data by automatically inserting hallucinations into training instances. Santhanam et al. (2021) and Honovich et al. (2021) construct new corpora for faithfulness classification in dialogue responses. They manually annotate the Wizard-of-Wikipedia dataset (Dinan et al., 2019), a knowledge grounded dialog dataset, by judging whether each response is hallucinated.
Faithfulness specific datasets can be better than NLI datasets because entailment or neutral labels of NLI datasets and faithfulness are not equivalent. For example, the hypothesis “Putin is U.S. president” can be considered to be either neutral to or entailed from the premise “Putin is president”. However, from the faithfulness perspective, the hypothesis contains unsupported information “U.S.”, which is deemed to be hallucination.
4.2.5. LM-based Metrics
These metrics leverage two language models (LMs) to determine if each token is supported or not: An unconditional LM is only trained on the targets (ground-truth references) in the dataset, while a conditional language model $LM_{x}$ is trained on both source and target data. It is assumed that the next token is inconsistent with the input if unconditional LM gets a smaller loss than conditional $LM_{x}$ during forced-path decoding (Filippova, 2020; Tian et al., 2020). We classify the generated token as hallucinatory if the loss from LM is lower. The ratio of hallucinated tokens to the total number of target tokens $|y|$ can reflect the hallucination degree.
4.3. Human Evaluation
Due to the challenging and imperfect nature of the current automatic evaluation of hallucinations in NLG, human evaluation (Shuster et al., 2021; Santhanam et al., 2021) is still one of the most commonly used approaches. There are two main forms of human evaluation: (1) scoring, where human annotators rate the hallucination level in a range; and (2) comparing, where human annotators compare the output texts with baselines or ground-truth references (Sun, 2010).
Multiple terminologies, such as faithfulness (Maynez et al., 2020; Cao et al., 2018; Chen et al., 2021b; Rashkin et al., 2021; Parikh et al., 2020; Filippova, 2020; Rashkin et al., 2021; Tian et al., 2020; Su et al., 2021; Xiao and Wang, 2021; Zhou et al., 2021b), factual consistency (Cao et al., 2020; Cao and Wang, 2021; Shen et al., 2021; Wu et al., 2021; Santhanam et al., 2021; Chen et al., 2020), fidelity (Chen et al., 2021a), factualness555uses the source input as the “fact”.  (Rebuffel et al., 2022), factuality4 (Dong et al., 2020), or on the other hand, hallucination (Huang et al., 2020a; Shuster et al., 2021; Santhanam et al., 2021; Dziri et al., 2021a; Liu et al., 2021b), fact contradicting (Nie et al., 2018) are used in the human evaluation of hallucination to rate whether the generated text is in accord with the source input. Chen et al. (2021b); Nie et al. (2019) use finer-grained metrics for intrinsic hallucination and extrinsic hallucination separately. Moreover, there are some broad metrics, such as Correctness (Balakrishnan et al., 2019; Bi et al., 2019; Wang et al., 2021a; Li et al., 2018), Accuracy (Yin et al., 2016; Li et al., 2021a), and Informativeness (Li et al., 2020c) considering both missing and additional contents (extrinsic hallucinations) compared to the input source.
5. Hallucination Mitigation Methods
Common mitigation methods can be divided into two categories, in accordance with two main contributors of hallucinations: Data-Related Methods, and Modeling and Inference Methods.
5.1. Data-Related Methods
5.1.1. Building a Faithful Dataset
Considering that noisy data encourage hallucinations, constructing faithful datasets manually is an intuitive method, and there are various ways to build such datasets: One way is employing annotators to write clean and faithful targets from scratch given the source (Gardent et al., 2017; Wen et al., 2015b), which may lack diversity (Parikh et al., 2020; Gururangan et al., 2018; Poliak et al., 2018). Another way is employing annotators to rewrite real sentences on the web (Parikh et al., 2020), or targets in the existing dataset (Wang, 2019). Basically, the revision strategy consists of three stages: (1) phrase trimming: removing phrases unsupported by the source in the exemplar sentence; (2) decontextualization: resolving co-references and deleting phrases dependent on context; (3) syntax modification: making the purified sentences flow smoothly. Meanwhile, other works (Honovich et al., 2021; Gabriel et al., 2021) leverage the model to generate data and instruct annotators to label whether these outputs contain hallucinations or not. While this approach is typically used to build diagnostic evaluation datasets, it has the potential to build faithful datasets.
5.1.2. Cleaning Data Automatically
In order to alleviate semantic noise issues, another approach is to find information that is irrelevant or contradictory to the input from the existing parallel corpus and then filter or correct the data. This approach is suitable for the case where there is a low or moderate level of noise in the original data (Filippova, 2020; Nie et al., 2019).
Some works (Liu et al., 2021b; Shen et al., 2021; Raunak et al., 2021) have dealt with the hallucination issue at the instance level by using a score for each source-reference pair and filtering out hallucinated ones. This corpus filtering method consists of several steps: (1) measuring the quality of the training samples in terms of hallucination utilizing the metrics described above; (2) ranking these hallucination scores in descending order; (3) selecting and filtering out the untrustworthy samples at the bottom. Instance-level scores can lead to a signal loss because divergences occur at the word level; i.e., parts of the target sentence are loyal to the source input, while others diverge (Rebuffel et al., 2022).
Considering this issue, other works (Nie et al., 2019; Dušek et al., 2019) correct paired training samples, specifically the input data, according to the references. This method is mainly applied in the data-to-text task because structured data are easier to correcte than utterances. This method consists of two steps: (1) utilizing a model to parse the meaning representation (MR), such as attribute-value pairs, from original human textual references; (2) using the MR extracted from the reference to correct the input MR through slot matching. This method will enhance the semantic consistency between input and output without abandoning a part of the dataset.
5.1.3. Information Augmentation
It is intuitive that augmenting the inputs with external information will obtain a better representation of the source. Because the external knowledge, explicit alignment, extra training data, etc., can improve the correlation between the source and target and help the model learn better task-related features. Consequently, a better semantic understanding helps alleviate the divergence from the source issue. Examples of the augmented information include entity information (Liu et al., 2021b), extracted relation triples from source document (Huang et al., 2020a; Cao et al., 2018) obtained by Fact Description Extraction, pre-executed operation results (Nie et al., 2018), synthetic data generated through replacement or perturbation  (Chen et al., 2021b; Lee et al., 2019), retrieved external knowledge (Shuster et al., 2021; Zhu et al., 2021; Gunel et al., 2019; Bi et al., 2019; Fan et al., 2019a), and retrieved similar training samples (Białecki et al., 2012).
These methods enforce a stronger alignment between inputs and outputs. However, they will bring challenges due to the gap between the original source and augmented information, such as the semantic gap between an ambiguous utterance and a distinct MR of structured data, and the format discrepancy between the structured knowledge graph and natural language.
5.2. Modeling and Inference Methods
5.2.1. Architecture
Encoder
The encoder learns to encode a variable-length sequence from input text into a fixed-length vector representation. As we mentioned above in Section 5.1.3, hallucination appears when the models lack semantic interpretation over the input. Some works have modified the encoder architecture in order to make it more compatible with input and learn a better representation. For example, Huang et al. (2020a) and Cao et al. (2018) propose a dual encoder, consisting of a sequential document encoder and a structured graph encoder to deal with the additional knowledge.
Attention
The attention mechanism is an integral component in neural networks that selectively concentrates on some parts of sequences while ignoring others based on dependencies (Vaswani et al., 2017). In order to encourage the generator to pay more attention to the source, Aralikatte et al. (2021) introduce a short circuit from the input document to the vocabulary distribution via source-conditioned bias. Krishna et al. (2021) employ sparse attention to improve the model‘s long-range dependencies in the hope of modeling more retrieved documents so as to mitigate the hallucination in the answer. Wu et al. (2021) adopt inductive attention, which removes potentially uninformative attention links by injecting pre-established structural information to avoid hallucinations.
Decoder
The decoder is responsible for generating the final output in natural language given input representations (Vaswani et al., 2017). Several work modified the decoder structures to mitigate hallucination, such as the multi-branch decoder (Rebuffel et al., 2022), uncertainty-aware decoder (Xiao and Wang, 2021), dual decoder, consisting of a sequential decoder and a tree-based decoder (Song et al., 2020a), and constrained decoder with lexical or structural limitations (Balakrishnan et al., 2019). Based on the observation that the “randomness” from sampling-based decoding, especially near the end of sentences, can lead to hallucination, (Lee et al., 2022) propose to iteratively reduce the “randomness” through time. These decoders improve the possibility of faithful tokens while reducing the possibility of hallucinatory ones during inference by figuring out the implicit discrepancy and dependency between tokens or restricted by explicit constraints. Since such decoders may have more difficulty generating fluent or diverse text, there is a balance to be struck between them.
5.2.2. Training
Figure 1. The frameworks of training methods.
Planning/Sketching
Planning is a common method to control and restrict what the model generates by informing the content and its order (Puduppully et al., 2019). Planning can be a separate step in a two-step generator (Liu et al., 2021b; Su et al., 2021; Chen et al., 2021b; Wang et al., 2021a; Puduppully and Lapata, 2021), which is prone to progressive amplification of the hallucination problem. Or be injected into the end-to-end model during generation (Xu et al., 2021a). Sketching has a similar function to planning, and can also be adopted for handling hallucinations (Wang et al., 2021a). The difference is that the skeleton is treated as a part of the final generated text. While providing more controllability, such methods also need to strike a balance between faithfulness and diversity.
Reinforcement Learning (RL)
As pointed out by Ranzato et al. (2016), word-level maximum likelihood training leads to the problem of exposure bias. Some works (Huang et al., 2020a; Su et al., 2021; Kong et al., 2019; Mesgar et al., 2021; Li et al., 2020c) adopt RL to solve the hallucination problem, which utilizes different rewards to optimize the model. The purpose of RL is for the agent to learn an optimal policy that maximizes the reward that accumulates from the environment (Uc-Cetina et al., 2021).The reward function is critical to RL and, if properly designed, it can provide training signals that help the model accomplish its goal of hallucination reduction. For example, Li et al. (2020c) propose a slot consistency reward which is the cardinality of the difference between generated template and the slot-value pairs extracted from input dialogue act. Improving the slot consistency can help reduce the hallucination phenomenon of missing or misplacing slot values in generated templates. Mesgar et al. (2021) attain persona consistency sub-reward via an NLI model to reduce the hallucinations in personal facts. Huang et al. (2020a) use a combination of ROUGE and the multiple-choice cloze score as the reward function to improve the faithfulness of summarization outputs. The cloze score is similar to the QA-based metric, measuring how well a QA model can address the questions by reading the generated summary (as context), where the questions are automatically constructed from the reference summary. As the above examples show, some RL reward functions for mitigating hallucination are inspired by existing automatic evaluation metrics. Although RL is challenging to learn and converge due to the extremely large search space, this method has the potential to obtain the best policy for the task without an oracle.
Multi-task Learning
Multi-task learning is also utilized for handling hallucinations in different NLG tasks. In this training paradigm, a shared model is trained on multiple tasks simultaneously to learn the commonalities of the tasks. The hallucination problem may be derived from the reliance of the training process on a single dataset, leading to the fact that the model fails to learn the actual task features. By adding proper additional tasks along with the target task during training, the model can suffer less from the hallucination problem. For example, Weng et al. (2020) and Garg et al. (2019) incorporate a word alignment task into the translation model to improve the alignment accuracy between the input and output, and thus faithfulness. Li et al. (2018) combine an entailment task with abstractive summarization to encourage models to generate summaries entailed by and faithful to the source. Li et al. (2021a) incorporate rationale extraction and the answer generation, which allows more confident and correct answers and reduces the hallucination problem. The Multi-task approach has several advantages, such as data efficiency improvement, overfitting reduction, and fast learning. It is crucial to choose which tasks should be learned jointly, and learning multiple tasks simultaneously presents new challenges of design and optimization (Crawshaw, 2020).
Controllable Generation
Current works treat the hallucination level as a controllable attribute in order to remain the hallucination in outputs at a low level. Controllable generation techniques such as controlled re-sampling (Rashkin et al., 2021), control codes that can be provided manually (Rashkin et al., 2021; Filippova, 2020; Wu et al., 2021), or predicted automatically (Wu et al., 2021) are leveraged to improve faithfulness. This method may require some annotated datasets for training. Considering that hallucination is not necessarily harmful and may bring some benefits, controllable methods can be further adapted to change the degree of hallucination to meet the demands of different real-world applications.
Other general training methods such as regularization (Lee et al., 2019; Müller et al., 2020; Kang and Hashimoto, 2020b) and loss reconstruction (Wang et al., 2020b; Li et al., 2020a; Wang and Sennrich, 2020) have also been proposed to tackle the hallucination problem.
5.2.3. Post-Processing
Post-processing methods can correct hallucinations in the output, and this standalone task requires less training data. Especially for noisy datasets where a large proportion of the ground truth references suffer from hallucinations, modeling correction is a competitive choice to handle the hallucination problem (Chen et al., 2021b). Chen et al. (2021b); Dong et al. (2020); Cao et al. (2020), and Dziri et al. (2021a) follow a generate-then-refine strategy. While the post-processing correction step tends to result in ungrammatical texts, this method allows researchers to utilise SOTA models which perform best in respect of other attributes, such as fluency, and then correct the results specifically for faithfulness by using small amounts of automatically generated training data.
6. Future Directions
Many studies have been conducted to tackle the hallucination problem in NLG and its downstream tasks. As mentioned above, we have discussed common metrics and mitigation methods to advance research in these fields. From a broader perspective, we wish to point out open challenges and potential directions in regard to metric and mitigation method.
6.1. Future Directions in Metrics Design
Fine-grained Metrics
Most of the existing hallucination metrics measure intrinsic and extrinsic hallucinations together as a unified metric. However, it is common for a single generation to have both types and a number of hallucinatory sub-strings. Fine-grained metrics that can distinguish between the two types of hallucinations will provide richer insight to researchers.
In order to implement a fine-graded metric, the first step would be to identify the exact location of the hallucinatory sub-strings correctly. However, some metrics such as those that are QA-based cannot identify the individual hallucinatory sub-strings. Improvements in this aspect would help improve the quality and explainability of the metrics. The next step would be to categorize the detected hallucinatory sub-strings. The hallucinatory sub-string will be intrinsic if it is wrong or nonsensical, and extrinsic if it is non-existing in the source context. Future work that explores an automatic method of categorization would be beneficial.
Fact-Checking
The factual verification of extrinsic hallucinations requires fact-checking against world knowledge, which can be time consuming and laborious. Leveraging an automatic fact-checking system for extrinsic hallucination verification is, thus, other future work that requires attention. Fact-checking consists of the knowledge evidence selection and claim verification sub-tasks, and the following are the remaining challenges associated with each sub-task.
The main research problem associated with the evidence selection sub-task is how to retrieve evidence from the world knowledge. Most of the literature leverages Wikipedia as the knowledge source (Thorne et al., 2018; Yoneda et al., 2018; Lee et al., 2020), which is only a small part of world knowledge. Other literature attempts to use the whole web as the knowledge source  (Etzioni et al., 2008; Magdy and Wanas, 2010). However, this method leads to another research problem – “how to ensure the trustworthiness of the information we use from the web” (Ginsca et al., 2015). Source-level methods that leverages the meta-information of the web source (e.g., web traffic, PageRank or URL structure) have been proposed to deal with this trustworthiness issue (Popat et al., 2016; Baly et al., 2018; Popat et al., 2018). Addressing the aforementioned issues to allow evidence selection against world knowledge will be an important future research direction.
For the verification subtask, verification models perform relatively well if given correct evidence (Lee et al., [n. d.]). However, it has been shown that verification models are prone to adversarial attacks and are not robust to negation, numerical or comparative words (Thorne et al., 2019). Improving this weakness of verification models would also be crucial because the factuality of a sentence can easily be changed by small word changes (i.e., changes in negations, numbers, and entities).
Generalization
Although we can see that the source and output text of different tasks are in various forms, investigating their relationship and common ground and proposing general metrics to evaluate hallucinations are worth exploring. Task-agnostic metrics with cross-domain robustness could help the research community to build a unified benchmark. It is also important and meaningful to build open-source platforms to collaborate and standardize the evaluation metrics for NLG tasks.
Incorporation of Human Cognitive Perspective
A good automatic metric should correlate with human evaluation. Humans are sensitive to different types of information. For instance, proper nouns are usually more important than pronouns in the generated text. Mistakes concerning named entities are striking to human users, but automatic metrics treat them equally if not properly designed. In order to address this issue, new metrics should be designed from the human cognitive perspective. The human ability to recognize salient information and filter the rest is evident in scenarios where the most important facts need to be determined and assessed. For instance, when signing an agreement, a prospective employee naturally skims the document to look at the entries with numbers first. In this way, humans classify what they believe is crucial.
Automatic check-worthy detection has the potential to be applied to improve the correlation with human judgement. Implementing the automatic human-like judgment mentioned above can further mitigate hallucination and improve NLG systems.
6.2. Future Directions in Mitigation Methods
General and robust data pre-processing approaches
Since the data format varies between downstream tasks, there is still a gap for data processing methods between tasks, and currently, no universal method is effective for all NLG tasks (Li et al., 2021b). Data pre-processing might result in grammatical errors or semantic transformation between the original and processed data, which can negatively affect the performance of generation. Therefore, we believe that general and robust data pre-processing methods can help mitigate the hallucinations in NLG.
Hallucinations in numerals
Most existing mitigation methods do not focus on the hallucination of numerals. However, the correctness of numerals in generated text, such as date, quantities and scalars are important for readers (Zhao et al., 2020; Thawani et al., 2021; Zhang et al., 2020b). For example, given the source document “The optimal oxygen saturation ($SpO_{2}$) in adults with COVID-19 who are receiving supplemental oxygen is unknown. However, a target $SpO_{2}$ of 92% to 96% seems logical, considering that indirect evidence from patients without COVID-19 suggests that an $SpO_{2}$ of ¡92% or ¿96% may be harmful.  [^6] https://www.covid19treatmentguidelines.nih.gov/management/critical-care/oxygenation-and-ventilation/”, the summary “The target oxygen saturation range for patients with COVID-19 is 82–86%.” includes wrong numbers, which could be fatal. Currently, some works (Nie et al., 2019; Thawani et al., 2021; Zhang et al., 2020b) point out that using commonsense knowledge can help to gain better numeral representation. And Zhao et al. (2020) alleviate numeral hallucinations by re-ranking candidate-generated summaries based on the verification score of quantity entities. Therefore, we believe that explicitly modeling numerals to mitigate hallucinations is a potential direction.
Extrinsic Hallucination Mitigation
Though many works on mitigating hallucinations have been published, most do not distinguish between intrinsic and extrinsic hallucination. Moreover, the main research focus has been on dealing with intrinsic hallucination, while extrinsic hallucination has been somewhat overlooked as it is more challenging to reduce (Huang et al., 2021). Therefore, we believe it is worth exploring different mitigation methods for intrinsic and extrinsic hallucinations, and relevant methods in fact-checking can be potentially used for this purpose.
Hallucination in long text
Many tasks in NLG require the model to process long input texts, such as multi-document summarization and generative question answering. We think adopting existing approaches to a Longformer (Beltagy et al., 2020)-based model could help encode long inputs. Meanwhile, part of dialogue systems need to generate long output text, in which the latter part may contradict history generation. Therefore, reducing self-contradiction is also an important future direction.
Reasoning
Misunderstanding facts in the source context will lead to intrinsic hallucination and errors. To help models understand the facts correctly requires reasoning over the input table or text. Moreover, if the generated text can be reasoned backwards to the source, we can assume it is faithful. There are some reasoning works in the area of dialogue (Cui et al., 2020; Ghosal et al., 2021; Wang et al., 2021b), but few in reducing hallucinations. Moreover, tasks with quantities, such as logical table-to-text generation, require numerical reasoning. Therefore, adding reasoning ability to the hallucination mitigation methods is also an interesting future direction.
Controllability
Controllability means the ability of models to control the level of hallucination and strike a balance between faithfulness and diversity (Rohrbach et al., 2018; Dziri et al., 2021a). As mentioned in Section 3, it is acceptable for chit-chat models to generate a certain level of hallucinatory content as long as it is factual. Meanwhile, for the abstractive summarization task, there is no agreement in the research community about whether factual hallucinations are desirable or not (Maynez et al., 2020). Therefore, we believe controllability merits attention when exploring hallucination mitigation methods.
7. Hallucination in Abstractive Summarization
Abstractive summarization aims to extract essential information from source documents and to generate short, concise, and readable summaries (Yu et al., 2021). Neural networks have achieved remarkable results on abstractive summarization. However, Maynez et al. (2020) observe that neural abstractive summarization models are likely to generate hallucinatory content that is unfaithful to the source document. Falke et al. (2019) analyze three recent abstractive summarization systems and show that 25% of the summaries generated from state-of-the-art models have hallucinated content. In addition, Zhou et al. (2021b) mention that even if a summary contains a large amount of hallucinatory content, it can achieve a high ROUGE (Lin, 2004) score. This has encouraged researchers to actively devise ways to improve the evaluation of abstractive summarization, especially from the hallucination perspective.
In this section, we review the current progress in automatic evaluation and the mitigation of hallucination, and list the remaining challenges for future work. In addition, it is worth mentioning that researchers have used various terms to describe the hallucination phenomenon, such as faithfulness, factual errors, and factual consistency, and we will use the original terms from their papers in the remainder of this section.
7.1. Hallucination Definition in Abstractive Summarization
The definition of hallucination in abstractive summarization follows that in Section 2. Specifically, we adopt the definition from (Maynez et al., 2020): given a document and its abstractive summary, a summary is hallucinated if it has any spans not supported by the input document. Once again, intrinsic hallucination refers to output content that contradicts the source, while extrinsic hallucination refers to output content that the source cannot verify. For instance, in Table 1, given the input article shown in the caption, an example of intrinsic hallucination is “The Ebola vaccine was rejected by the FDA in 2019,” because this statement contradicts the given content “The first vaccine for Ebola was approved by the FDA in 2019 in the US”. And an example of extrinsic hallucination is “China has already started clinical trials of the COVID-19 vaccine,” because this statement is not mentioned in the given content. We can neither find evidence of it from the input article nor assert that it is wrong.
Pagnoni et al. (2021) define fine-grained types of factual errors in summaries. As mentioned in 2.3, since the “fact” here refers to source knowledge, “factual error” can be treated as hallucination, and we can adopt this classification as a sub-type of hallucination. They establish three categories as semantic frame error, discourse error, and content verifiability error.
7.2. Hallucination Metrics in Abstractive Summarization
Existing metrics for hallucination in abstractive summarization are mainly model-based. Following (Huang et al., 2021), we divide the hallucination metrics into two categories: (1) unsupervised metrics and (2) semi-supervised metrics. Note that existing hallucination metrics evaluate both intrinsic and extrinsic hallucinations together in one metric because it is difficult to automatically distinguish between them.
7.2.1. Unsupervised Metrics
Given that hallucination is a newly emerging problem, there are only a few hallucination-related datasets. Therefore, researchers have proposed to adopt other datasets to build unsupervised hallucination metrics. There are three types of such unsupervised metrics: (1) information extraction (IE)-based metrics, (2) natural language inferencing (NLI)-based metrics, (3) question answering (QA)-based metrics.
IE-based Metrics
As mentioned in Section 4, IE-based metrics leverage IE models to extract knowledge as relation tuples (subject, relation, object) from both the generation and knowledge source to analyze the factual accuracy of the generation (Goodrich et al., 2019). However, IE models are not 100% reliable yet (making errors in the identification of the relation tuples). Therefore, Nan et al. (2021) propose an entity-based metric relying on the Named-Entity Recognition model, which is relatively more robust. Their metric builds on the assumption that there will be a different set of named entities in the gold and generated summary if there exists hallucination.
NLI-based Metrics
As mentioned in Section 4, the NLI-model (textual entailment model) can be utilized to measure hallucination based on the assumption that a faithful summary will be entailed by the gold source. However, Falke et al. (2019) discover that models trained on NLI datasets can not transfer well to abstractive summarization tasks, degrading the reliability of NLI-based hallucination metrics. To improve NLI models for hallucination evaluation, they release collected annotations as additional test data. Other efforts have also been made to further improve NLI models. Mishra et al. (2021) find that the low performance of NLI-based metrics is mainly caused by the length of the premises in NLI datasets being shorter than the source documents in abstractive summarization. Thus, the authors propose to convert multiple-choice reading comprehension datasets into long premise NLI datasets automatically. The results indicate that long-premise NLI datasets help the model achieve a higher performance than the original NLI datasets. In addition, Laban et al. (2022) introduce a simple but efficient method called SUMMACConv by applying NLI models to sentence units that are segmented from documents. The performance of their model is better than applying NLI models to the whole document.
QA-based Metrics
QA-based metrics measure the knowledge overlap or consistency between summaries and the source documents based on the intuition that QA models will achieve similar answers if the summaries are factually consistent with the source documents. QA-based metrics such as FEQA (Durmus et al., 2020), QAGS (Wang et al., 2020a), and QuestEval (Scialom et al., 2021) follow three steps to obtain a final score: (1) a QG model generates questions from the summaries, (2) a QA model obtains answers from the source documents, and (3) calculate the score by comparing the set of answers from source documents and the set of answers from summaries. The results show that these reference-free metrics have substantially higher correlations with human judgments of faithfulness than the baseline metrics. Gabriel et al. (2021) further analyze the FEQA and find that the effectiveness of QA-based metrics depends on the question. They also provide a meta-evaluation framework that includes QA metrics.
7.2.2. Semi-Supervised Metrics
Semi-supervised metrics are trained on the synthetic data generated from summarization datasets. Trained on these task-specific corpora, models can judge whether the generated summaries are hallucinatory. Kryscinski et al. (2020) propose a weakly supervised model named FactCC for evaluating factual consistency. The model is trained jointly for three tasks: (1) checking whether the synthetic sentences remain factually consistent, (2) extracting supporting spans in the source documents, and (3) extracting inconsistent spans in the summaries, if any exist. They transfer this model to check whether the summaries generated from summarization models are factually consistent. Results show that the performance of their FactCC model surpasses the classifiers trained on the MNLI or FEVER datasets. Zhou et al. (2021b) introduce a method to fine-tune a pre-trained language model on synthetic data with automatically inserted hallucinations in order to detect the hallucinatory content in summaries. The model can classify whether spans in the machine-generated summaries are faithful to the article. This method shows higher correlations with human factual consistency evaluation than the baselines.
7.3. Hallucination Mitigation in Abstractive Summarization
Recently, many approaches have been proposed to reduce the hallucination phenomenon in abstractive summarization.
7.3.1. Architecture Method.
Seq-to-seq (Sutskever et al., 2014) models are widely used and achieve state-of-the-art performance in abstractive summarization. Researchers have made modifications to the architecture design of the seq-to-seq models to reduce hallucinated content in the summaries. We describe various efforts made to improve the encoder, decoder, or both the encoder and decoder of the seq-to-seq models.
Encoder
Zhu et al. (2021) propose to use an explicit graph neural network (GNN) to encode the fact tuples extracted from source documents. In addition to an explicit graph encoder, Huang et al. (2020a) further design a multiple-choice cloze test reward to encourage the model to better understand entity interactions. Moreover, Gunel et al. (2019) use external knowledge from Wikipedia to make knowledge embeddings, which the results show improve factual consistency.
Decoder
Song et al. (2020a) present the incorporation of a sequential decoder with a tree-based decoder to generate a summary sentence and its syntactic parse. This joint generation is performed improve faithfulness. Aralikatte et al. (2021) introduce the Focus Attention Mechanism, which encourages decoders to generate tokens similar or topical to the source documents. The results on the BBC extreme summarization task show that models augmented with the Focus Attention Mechanism generate more faithful summaries.
Encoder-decoder
Cao et al. (2018) extract fact descriptions from the source text and apply a dual-attention seq-to-seq framework to force the summaries to be conditioned on both source documents and the extracted fact descriptions. Li et al. (2018) propose an entailment-aware encoder and decoder with multi-task learning which incorporates the entailment knowledge into abstractive summarization models.
7.3.2. Training Method
Aside from architecture modification, some works improved the training approach to reduce hallucination. Cao and Wang (2021) introduce a contrastive learning method to train summarization models. The positive training data are reference summaries, while the negative training data are automatically generated hallucinatory summaries, and the contrastive learning system is trained to distinguish between them. In the dialogue summarization field, Tang et al. (2021) propose another contrastive fine-tuning strategy, named CONFIT, that can improve the factual consistency and overall quality of summaries.
7.3.3. Post-Processing Method
Some works carry out post-editing to reduce the hallucination of the model-generated summaries, which are viewed as draft summaries. Dong et al. (2020) propose SpanFact, a pair of factual correction models that use knowledge learned from QA models to correct the spans in the generated summaries. Similar to SpanFact, Cao et al. (2020) introduce a post-editing corrector module to identify and correct hallucinatory content in generated summaries. The corrector module is trained on synthetic data which are created by adding a series of heuristic transformations to reference summaries. Zhao et al. (2020) present HERMAN, a system that learns to recognize quantities (dates, amounts of money, etc.) in the generated summary and verify their factual consistency with the source text. According to the quantity hallucination score, the system chooses the most faithful summary where the source text supports its quantity terms from the candidate-generated summaries. Chen et al. (2021b) introduce a contrast candidate generation and selection system to do post-processing. The contrast candidate generation model replaces the named entities in the generated summaries with ones present in the source documents, and the contrast candidate selection model will select the best candidate as the final output summary.
7.4. Future Directions in Abstractive Summarization
Factual Hallucination Evaluation
Factual hallucinations contain information not found in source content, though it is factually correct. In the summarization task, this kind of hallucination could lead to better summaries. However, there is little work focused on evaluating factual hallucination explicitly. Fact-checking approaches could be potentially used in this regard.
Extrinsic Hallucination Mitigation
There has been little research on extrinsic hallucinations as it is more challenging to detect and mitigate content based on world knowledge. We believe it is worth exploring extrinsic hallucination in terms of evaluation metrics and mitigation methods.
Hallucination in Dialogue Summarization
In conversational data, the discourse relations between utterances and co-references between speakers are more complicated than from, say, news articles. For example, Zhong et al. (2021) show that 74% of samples in the QMSum dataset consist of inconsistent facts. We believe exploring the hallucination issue in dialogue summarization is an important and special component of research into hallucination in abstractive summarization.
8. Hallucination in Dialogue Generation
Dialogue generation is an NLG task that automatically generates responses according to user utterances. The generated responses are required to be fluent, coherent, and consistent with the dialogue history. The dialogue generation task can be divided into two sub-tasks: (1) task-oriented dialogue generation; (2) open-domain dialogue generation. A task-oriented dialogue system aims to complete a certain task according to a user query in a specific domain, such as restaurant booking, hotel recommendation, and calendar checking. Meanwhile, an open-domain dialogue system aims to establish a multi-turn, long-term conversation with users while providing the users with an engaging experience.
8.1. Hallucination Definition in Dialogue Generation
The hallucination problem also exists in the dialogue generation task. It is important to note that a dialogue system is expected either to provide the user with the required information or to provide an engaging response without repeating utterances from the dialogue history. Thus, the tolerance for producing proper “hallucination” from the dialogue history is relatively higher.
The definition of hallucination in this task can be adopted from the general definition as follows: (1) Intrinsic hallucination: the generated response is contradictory to the dialogue history or the external knowledge sentences. In the examples of intrinsic hallucination shown in Table 1, we can verify that the output contradicts the inputs: In one example, the input is a “moderate” price range, but the model mistakenly generates a sentence with a “high” price range. In another case, the confusion of the names “Roger Federer” and “Rafael Nadal” causes the output generation of “Roger Nadal”. (2) Extrinsic hallucination: the generated response is hard to verify with the dialogue history or the external knowledge sentences. Responses with extrinsic hallucination are impossible to verify with the given inputs. “Pickwick hotel” might be “in san diego”, and Djokovic may have been “in the top ten singles players of the world”. However, we do not have enough information to check the truth of these statements.
In the following sections, the hallucination problem in open-domain and task-oriented dialogue generation tasks will be separately discussed according to the their natures.
8.2. Open-domain Dialogue Generation
While the term “hallucination” seems to have newly emerged in the NLP field, a related behavior, “inconsistency”, of neural models has been widely discussed. This behavior has been pointed out as a shortcoming of generation-based approaches for open-domain chatbots (Huang et al., 2020b; Ma et al., 2020; Roller et al., 2020). Two possible types of inconsistency occur in open-domain dialogue generation: (1) inconsistency among the system utterances, such as when the system contradicts its previous utterance; (2) inconsistency with an external source, such as factually incorrect utterances. Whereas the first type is described using the term ”inconsistency” (Welleck et al., 2019b; Li et al., 2020b; Zhang et al., 2021a) or ”incoherence” (Dziri et al., 2019; Beyer et al., 2021), some have recently started to call the second type ”hallucination” (Roller et al., 2021; Mielke et al., 2020). Self-inconsistency can be considered as an intrinsic hallucination problem, while the external inconsistency involves both intrinsic and extrinsic hallucinations, depending on the reference source.
As mentioned earlier, a certain level of hallucination may be acceptable in open-domain chit-chat as long as it does not involve severe factual issues. Moreoever, it is almost impossible to verify factual correctness since the system usually lacks a connection to external resources. With the introduction of knowledge-grounded dialogue tasks (Zhou et al., 2018; Dinan et al., 2019), which provide an external reference, however, there has been more active discussion of hallucination in open-domain dialogue generation.
8.2.1. Self-Consistency
In end-to-end generative open-domain dialogue systems, the inconsistency among system utterances has been pointed out as the bottleneck to human-level performance  (Vinyals and Le, 2015). We often observe an inconsistency in the answers to semantically similar yet not identical questions. For example, a system may answer the questions of “What is your name?” and “May I ask your name?” with different responses. Persona consistency has been the center of attention (Li et al., 2016; Zhang et al., 2018) and it is one of the most obvious cases of self-contradiction regarding the character of the dialogue system. ”Persona” is defined as the character that a dialogue system plays during a conversation, and can be composed of identity, language behavior, and an interaction style (Li et al., 2016). While some works has set their objective as teaching models to utilize speaker-level embeddings (Li et al., 2016; Madotto et al., 2019), others condition generation with a set of descriptions about a persona, which we will discuss in detail in the next section.
8.2.2. External Consistency
Besides self-consistency, an open-domain dialogue system should also generate persona-consistent and informative responses corresponding so as to user utterances to further engage with the user during conversation. In this process, an external resource containing explicit persona information or world knowledge is introduced into the system to assist the model generation process.
The PersonaChat datasets (Zhang et al., 2018; Dinan et al., 2020) have accelerated research into persona consistency (Hancock et al., 2019; Kulikov et al., 2019; Mazaré et al., 2018; Yavuz et al., 2019; Zemlyanskiy and Sha, 2018; Wolf et al., 2019; Zhang et al., 2020c). In PersonaChat datasets, each conversation has persona descriptions such as “I like to ski” or “I am a high school teacher” attached. By conditioning the response generation on the persona description, a chit-chat model is expected to acquire an ability to generate a more persona-consistent response. Lately, the application of NLI methods (Li et al., 2020b; Song et al., 2020b) or reinforcement learning frameworks (Mesgar et al., 2021) have been investigated. Although these methods conditioned on the PersonaChat datasets have been successful, further investigation of approaches that do not rely on a given set of persona descriptions is necessary because such descriptions are not always available, and covering every aspect of a persona with them is impossible.
In addition to PersonaChat-related research, the knowledge-grounded dialogue (KGD) task in the open-domain requires the model to generate informative responses with the help of an external knowledge graph (KG) or knowledge corpus (Zhou et al., 2018; Dinan et al., 2019). Hallucination in conversations, which is also considered as a factual consistency problem, has raised much research interest recently (Dziri et al., 2021a; Shuster et al., 2021; Rashkin et al., 2021; Santhanam et al., 2021). Here, we continue to split the hallucination problem in the KGD task into intrinsic hallucination and extrinsic hallucination. Most of the KGD works tackle the hallucination problem when responses contain information that contradicts (intrinsic) or cannot be found in the provided knowledge input (extrinsic). Since world knowledge is enormous and ever-changing, the extrinsic hallucination may be factual but hard to verify. Dziri et al. (2021a) further adopt the same definition of hallucination as mentioned above to the knowledge graph-grounded dialogue task, where intrinsic hallucination indicates the case of misusing either the subject or object of the knowledge triple; and extrinsic hallucination indicates that there is no corresponding valid knowledge triple in the gold reference knowledge. Recently, there have been some attempts to generate informative responses without explicit knowledge inputs, but with the help of the implicit knowledge inside large pre-trained language models instead (Xu et al., 2021b; Zhou et al., 2021a) during the inference time. Under this setting, the study of extrinsic hallucination is of great value but still poorly investigated.
8.2.3. Hallucination Metrics
For generation-based dialogue systems, especially open-domain chatbots, the hallucination evaluation method remains an open problem (Roller et al., 2020). As of now, there is no standard metric. Therefore, chatbots are usually evaluated by humans on factual consistency or factual correctness (Wu et al., 2021; Santhanam et al., 2021). We also introduce some automatic statistical and model-based metrics as a reference, which will be described in more detail below.
Variants of F1 Metrics
Knowledge F1 (KF1) measures the overlap between the generated responses and the gold knowledge sentences to which the human referred for conversation during dataset collection (Shuster et al., 2021). KF1 attempts to capture whether a model can generate knowledgable responses by correctly utilizing the relevant knowledge. KF1 is only available for datasets with labeled ground-truth knowledge. Shuster et al. (2021) further propose Rare F1 (RF1), which only considers the infrequent words in the dataset when calculating F1 to avoid influence from the common uni-grams. The authors define an infrequent word as being in the lower half of the cumulative frequency distribution of the reference corpus.
Model-based Metric
Natural language has its natural on the flexibility of the surface forms with the same semantics, so overlap-based metrics cannot provide the comprehensive evaluation. Recently, several works have proposed evaluation metrics for measuring consistency, such as using natural language inference (NLI) (Welleck et al., 2019b; Dziri et al., 2019), training learnable evaluation metrics (Zhang et al., 2021a), or releasing an additional test set for coherence (Beyer et al., 2021). These methods are more flexible and supports the generated responses with different surface forms. For the KGD task, Dziri et al. (2021b) propose the BEGIN benchmark, which consists of samples taken from Dinan et al. (2019) with additional human annotation and a new classification task extending the NLI paradigm. Honovich et al. (2021) present a trainable metric for the KGD task, which also applies NLI. It is also noteworthy that Gupta et al. (2021) propose datasets that can benefit fact-checking systems specialized for dialogue systems. The Conv-FEVER corpus (Santhanam et al., 2021) is a factual consistency detection dataset, which was created by adapting the Wizard-of-Wikipedia dataset (Dinan et al., 2019). It consists of both factually consistent and inconsistent responses and can be used to train a classifier to detect factually inconsistent responses with respect to the knowledge provided.
8.2.4. Mitigation Methods
The hallucination issue can be mitigated by data pre-processing, which includes introducing extra information into the data. Shen et al. (2021) propose a measurement based on seven attributes of the dialogue quality, including self-consistency. Based on this measurement, the untrustworthy samples which get lower scores are filtered out from the training set to improve the model performance in terms of self-consistency (i.e., intrinsic hallucination). Shuster et al. (2021) conduct a comprehensive investigation on a retrieval-augmented KGD task where a retriever is introduced to the system for knowledge selection. The authors study several key problems, such as whether retrieval helps reduce hallucinations and how the generation should be augmented with the retrieved knowledge. The experimental results show that retrieval helps substantially in improving performance on KGD tasks and in reducing the hallucination in conversations without sacrificing conversational ability.
Rashkin et al. (2021) introduce a set of control codes and concatenate them with dialogue inputs to reduce the hallucination by forcing the model to be more aware of how the response relies on the knowledge evidence in the response generation. Some researchers have also tried to reduce hallucinated responses during generation by improving dialogue modeling. Wu et al. (2021) apply inductive attention into transformer-based dialogue models, and potentially uninformative attention links are removed with respect to a piece of pre-established structural information between the dialogue context and the provided knowledge. Instead of improving the dialogue response generation model itself, Dziri et al. (2021a) present a response refinement strategy with a token-level hallucination critic and entity-mention retriever, so that the original dialogue model is left without retraining. The former module is designed to label the hallucinated entity mentioned in the generated responses, while the retriever is trained to retrieve more faithful entities from the provided knowledge graph. RHO (ji-etal-2023-rho) is a framework that uses three mechanisms to tackle hallucinations, namely, local knowledge grounding, global knowledge grounding, and response re-ranking to tackle hallucinations in open-domain dialogues, and has been empirically shown to perform this.
8.3. Task-oriented Dialogue Generation
A task-oriented dialogue system is often composed of several modules: a natural language understanding (NLU) module, a dialogue manager (DM), and a natural language generation (NLG) module (Gao et al., 2018; Jurafsky and Marin, 2019). Intrinsic hallucination can occur between the DM and NLG, where a dialogue act such as recommend(NAME=peninsula hotel, AREA=tsim sha tsui) is transformed into a natural language representation “the hotel named peninsula hotel is located in tsim sha tsui area.” (Balakrishnan et al., 2019; Li et al., 2020c).
8.3.1. Hallucination Metrics
To evaluate hallucination, Li et al. (2020c) and Balakrishnan et al. (2019) combine traditional metrics such as the BLEU score and human evaluation as well as hallucination-specific automatic metrics. Following previous works such as (Wen et al., 2015a; Dušek and Jurčíček, 2016), and (Tran and Nguyen, 2017), Li et al. (2020c) use the slot error rate, which is computed by $(p+q)/N$, where $N$ represents the total number of slots extracted by another model in the dialogue act. Here, $p$ stands for the missing slots in the generated template, and $q$ is the number of redundant slots. On the other hand, Balakrishnan et al. (2019) introduce a novel metric called the tree accuracy, which determines whether the prediction’s tree structure is identical to that of the input meaning representations.
8.3.2. Mitigation Methods
While Balakrishnan et al. (2019) propose to adopt tree-structured semantic representations and add constraints on decoding, Li et al. (2020c) frame a reinforcement learning problem to which they apply a bootstrapping algorithm to sample training instances and then leverage a reward related to slot consistency. Recently, there has emerged another line of research in task-oriented dialogue, which is to build a single end-to-end system rather than connecting several modules (e.g., Eric and Manning (2017); wu2018globaltolocal; Madotto et al. (2018, 2020a)). As discussed in previous sections of this paper, there is potential for such end-to-end systems to produce extrinsic hallucinations, yet this remains less explored. For example, a model might generate a response with an entity that appears out of nowhere. In the example of hotel recommendation in Hong Kong given above, a model could generate a response such as “the hotel named raffles hotel is located in central area, [^7] Raffles Hotel is a hotel located in Downtown Core, Singapore.” which cannot be verified from the knowledge base of the system.
8.4. Future Directions in Dialogue Generation
Self-Contradiction in Dialogue Systems
One of the possible reasons for self-contradiction is that current dialogue systems tend to have a short memory of dialogue history (Roller et al., 2020). Firstly, common dialogue datasets provide several turns of conversation, yet these are not long enough to assess a model’s ability to deal with a long context. To overcome this, Xu et al. (2021c) introduce a new dataset that consists of, on average, over 40 utterances per episode. Secondly, we often truncate dialogue history into fewer turns to fit into models such as Transformer-based architectures, which makes it difficult for a model to memorize the past. In addition to the works on dialogue summarization, e.g., Gliwa et al. (2019), it would be beneficial to apply other works which are aiming to grasp the longer context but do not focus on dialogue generation (Beltagy et al., 2020; Zaheer et al., 2020; Zhao et al., 2021).
Fact-checking in dialogue systems
In addition to the factual consistency in responses from knowledge grounded dialogue systems, fact-checking is a future direction in dealing with the hallucination problem in dialogue systems (Gupta et al., 2021). Dialogue fact-checking involves verifiable claim detection, which is an important line in distinguishing hallucination-prone dialogue, and evidence retrieval from an external source. This fact-checking in the dialogue system could be utilized not only as an evaluation metric for facilitating factual consistency but also to model such a system.
9. Hallucination in Generative Question Answering
Generative question answering (GQA) aims to generate an abstractive answer rather than extract an answer to a given question from provided passages (Fan et al., 2019b; Li et al., 2021a). It is an important task since many of the everyday questions that humans deal with and pose to search engines require in-depth explanations (Khashabi et al., 2021) (e.g., why/how..?), and the answers are normally long and cannot be directly extracted from existing phrase spans. A GQA system can be integrated with a search engine (Metzler et al., 2021) to empower more intelligent search or combined with a virtual conversation agent to enhance user experience.
Normally, a GQA system involves searching an external knowledge source for information relevant to the question. Then it generates the answer based on the retrieved information (Krishna et al., 2021). In most cases, no single source (document) contains the answer, and multiple retrieved documents will be considered for answer generation. Those documents may contain redundant, complementary, or contradictory information. Thus, hallucination is common in the generated answers.
The hallucination problem is one of the most important challenges in GQA. Since an essential goal of a GQA system is to provide factualy-correct answers given the question, hallucination in the answer will mislead the user and damage the system performance dramatically.
9.1. Hallucination Definition in GQA
As a challenging yet under-explored task, there is no standard definition of hallucination in GQA. However, almost all the works on GQA (Fan et al., 2019b; Krishna et al., 2021; Nakano et al., 2021; Su et al., 2022) involve a human evaluation process, in which the factual correctness measuring the faithfulness of the generated answer can be seen as a measurement of the hallucination; i.e., the more faithful the answer is, the less hallucinated content it contains. The most recent such work  (Li et al., 2021a) uses the term semantic drift, which indicates how the answer drifts away from a correct one during generation, and this can also be seen as a specific definition of hallucination in GQA.
In line with the general categorization of hallucination in Section  2.1, we give two concrete hallucination examples in GQA in Table  1. The sources of both questions are Wikipedia web pages. For the first question, “dow jones industrial average please?”, the generated answer “index of 30 major U.S. stock indexes” conflicts with the statement “of 30 prominent companies listed on stock exchanges in the United States” from Wikipedia. So we categorize it as an intrinsic hallucination. For the second example, the sentences “The definition of a Sadducee is a person who acts in a deceitful or duplicitous manner. An example of a Sadduceee is a politician who acts deceitfully in order to gain political power” in the generated answer can not be verified from the source documents; thus, we categorize it as an extrinsic hallucination.
9.2. Hallucination-related Metrics in GQA
Currently, there is no automatic metric to evaluate hallucination in QGA specifically. While most works on GQA use automatic evaluation metrics such as ROUGE score and F1 to measure the quality of the answer, these N-gram overlap-based metrics are not a meaningful way to evaluate hallucination due to their poor correlation with human judgments, as indicated by Krishna et al. (2021). On the other hand, almost all the GQA-related work involves a human evaluation process as a complement to the automatic evaluation. Normally, human annotators will be asked to assign a score indicating the faithfulness of the answer, which can also be viewed as a measurement of the answer hallucination. However, the metrics obtained via human evaluation come normally from a small sample of the data.
Metrics such as semantic overlap  (Sellam et al., 2020), a learned evaluation metric based on BERT that models human judgments, could be considered a better measurement of hallucination for GQA. Other metrics such as the factual correctness can also be considered as a way to measure hallucination in GQA. Zhang et al. (2020a) propose to explicitly measure the factual correctness of a generated text against the reference by first extracting facts via an information extraction (IE) module. Then they define and measure the factual accuracy score to be the ratio of facts in the generation text equal to the corresponding facts in the reference.
Factual consistency, which measures the faithfulness of the generated answer given its source documents, can be employed as another way to measure hallucination in GQA. Durmus et al. (2020); Wang et al. (2020a) propose an automatic QA-based metric to measure faithfulness in summary, leveraging the recent advances in machine reading comprehension. They first use a question generation model to construct question-answer pairs from the summary, and then a QA model is applied to extract short answer spans from the given source document for the question. The extracted answers that do not match the provided answers indicate unfaithful information in the summary. While these metrics were first proposed in summarization works, they can be easily adopted in generative QA to measure hallucinations in the generated long-form answer.
The most recent work on GQA by  Su et al. (2022) proposed to estimate the faithfulness of the generated long-form answer via zero-shot short answer recall on extractive QA datasets. They first generate long-form answers for questions from two extractive QA datasets Natural Questions(NQ) (Kwiatkowski et al., 2019) and HotpotQA (Yang et al., 2018), both of which contains large-scale question-answer pairs, then they measure the ratio of golden short answer span contained in the generated long answer as an estimation of faithfulness of the generated long-answer. While the idea is similar to the factual consistency metric in summarization work (Durmus et al., 2020), and also matches with our intuition to some extent, its correlation with human evaluation on faithfulness has not been verified.
9.3. Hallucination Mitigation in GQA
Unlike conditional text generation tasks such as summarization, or data-to-text generation, in which the source documents are provided and normally related to the target generation, the hallucination problem in GQA is more complicated. Generally speaking, it might come from two sources: 1) the incompetency of the retriever, which retrieves documents irrelevant to the answer, and 2) the intrinsic and extrinsic hallucination in the conditional generation model itself. Normally these two parts are interconnected and cause hallucinations in the answer.
Early works on GQA mostly tried to improve the faithfulness of the answer by investigating reliable external knowledge sources or incorporating multiple information sources. Yin et al. (2016) propose Neural Generative Question Answering (GENQA), an end-to-end model that generates answers to simple factoid questions based on the knowledge base, while Bi et al. (2019) propose the Knowledge-Enriched Answer Generator (KEAG) to generate a natural answer by integrating facts from four different information sources, namely, questions, passages, vocabulary, and knowledge. Nevertheless, these methods rely on the existence of high-quality, relevant resources which are not easily available.
Recent works focus more on the conditional generation model. Fan et al. (2019a) construct a local knowledge graph for each question to compress the information and reduce redundancy from the retrieved documents, which can be viewed as an early trial to mitigate hallucination. Li et al. (2021a) propose Rationale-Enriched Answer Generator (REAG), in which they add an extraction task to obtain the rationale for an answer at the encoding stage, and the decoder is expected to generate the answer based on both the extracted rationale and original input. The recent work (Krishna et al., 2021) employs a Routing Transformer (RT), a sparse attention-based Transformer-based model that employs local attention and mini-batch k-means clustering for long-range dependence, as the answer generator in the hope of modeling more retrieved documents to mitigate the hallucination in the answer. Su et al. (2022) propose a framework named RBG (read before generate), to jointly models answer generation with machine reading. They augment the generation model with fine-grained, answer-related salient information predicted by the MRC module, to enhance answer faithfulness. Such methods can exploit and utilize the information in the original input better, while they require the extra effort of building models to extract that information.
Most recently, Lin et al. (2021) propose a benchmark, which comprises 817 questions that span 38 categories, to measure the truthfulness of a language model in the QA task. This work investigates the performances of GPT-3 (Brown et al., 2020), GPT-Neo/J (Wang and Komatsuzaki, 2021), GPT-2 (Radford et al., 2019) and a T5-based model (Raffel et al., 2020). The results suggest that simply scaling up the model is less promising than fine-tuning it in terms of improving truthfulness since larger models are better at learning the training distribution from web data and thus tend to produce more imitative falsehoods. In another recent work, Nakano et al. (2021) fine-tune GPT-3 to answer long-form questions with a web-browsing environment, which allows the model to navigate the web as well as use human feedback to optimize answer quality directly using imitation learning (Hussein et al., 2017). While this method seems promising, it also hinges on how that feedback is processed.
9.4. Future Directions in GQA
While GQA is challenging yet under-explored, many possible directions could be explored to improve the answer quality and mitigate hallucination. First, better automatic evaluation metrics are needed to measure hallucination. The previously mentioned metrics, such as the semantic overlap between the generated answer and the ground-truth answer, the faithfulness of the generated answer, and factual consistency between the answer and the source documents, only consider one aspect of hallucination. Metrics that can consider all the factors related to hallucination (such as semantic overlap, faithfulness, or factual consistency) could be designed. Second, datasets with hallucination annotations should be proposed since none of the current GQA datasets have that information. Another possible direction to mitigate hallucination in the answer is improving the performance of the models. We need better retrieval models that retrieve relevant information according to queries and generation models that can synthesize more accurate answers from multi-source documents.
10. Hallucination in Data-to-Text Generation
Data-to-Text Generation is the task of generating natural language descriptions conditioned on structured data (Kukich, 1983; McKeown, 1992), such as tables (Parikh et al., 2020; Wiseman et al., 2017), database records (Chisholm et al., 2017), and knowledge graphs (Gardent et al., 2017). Although this field has been recently boosted by neural text generation models, it is well known that these models are prone to hallucinations (Wiseman et al., 2017) because of the gap between structured data and text, which may cause semantic misunderstanding and erroneous correlation. Moreover, the tolerance of hallucination is very low when this task is applied to the real world, such as in the case of patient information table description (Thomson and Reiter, 2020), and analysis of experimental results tables in a scientific report. Recent years have seen a growth of interest in hallucinations in Data-to-Text Generation, and researchers have proposed works from the aspect of evaluation and mitigation.
10.1. Hallucination Definition in Data-to-Text Generation
The definition and categories of hallucination in Data-to-Text Generation follow the descriptions in Section 2. We follow the general hallucination definition in this task: (1) Intrinsic Hallucinations: the generated text contains information that is contradicted by the input data (Nie et al., 2019). For example, in Table 1, “The Houston Rockets (18-4)” uses the information “[TEAM: Rockets, CITY:Houston, WIN:18, LOSS: 5]” in the source table. However, “(18-4)” is contradicted by “[LOSS: 5]” and it should be “(18-5)”. (2) Extrinsic Hallucinations: the generated text contains extra information irrelevant to the input  (Dhingra et al., 2019; Nie et al., 2019). For example, in Table 1, “Houston has won two straight games and six of their last seven.” is not mentioned in the source table (Wang, 2019).
10.2. Hallucination Metrics in Data-to-Text Generation
Statistical
PARENT (Dhingra et al., 2019) measures the accuracy of table-to-text generation by aligning n-grams from the reference description $R$ and generated texts $G$ to the table $T$. And it is the average F-score by combining the entailment precision and recall. Wang et al. (2020b) modify PARENT and denote this table-focused version as PARENT-T. Different from PARENT, which evaluates i-th instance $(T_{i},R_{i},G_{i})$, PARENT-T ignores the reference description R and evaluates each instance $(T_{i},G_{i})$.
IE-based
Liu et al. (2021b) estimate hallucination with two entity-centric metrics: table record coverage (the ratio of covered records in a table) and hallucinated ratio (the ratio of hallucinated entities in text). This metric firstly uses entity recognition to extract the entities of input and generated output, then aligns these entities by heuristic matching strategies, and finally calculates the ratios of faithful and hallucinated entities separately. Moreover, there are some general post-hoc IE-based metrics that could be applied to hallucination evaluation, such as Slot Error Rate (SER) (Xu et al., 2021a), Content Selection (CS), Relation Generation (RG), and Content Ordering (CO) (Wiseman et al., 2017; Wang, 2019).
QA-based
Data-QuestEval (Rebuffel et al., 2021) adapt QuestEval (Scialom et al., 2021) from summarization into data-to-text generation. First, a textual QG model is trained on a textual QA dataset. For each sample (structured data, textual descriptions), the textual QG model generates synthetic problems based on the descriptions. The structured data, textual descriptions (answers), and synthetic questions make up a synthetic QG/QA dataset to train synthetic QA/QG models. Then, the synthetic QG model generates questions based on the textual description to be evaluated. The synthetic QA model then generates answers based on a synthetic question and the structured input data. Finally, BERTScore (Zhang et al., 2019a) measures the similarity between the generated answer and description, indicating faithfulness.
NLI-based
Dušek and Kasner (2020) recognize the textual entailment between the input data and the output text for both omissions and hallucinations with an NLI model. This work measures the semantic accuracy in two directions: check omissions by inferring whether the input fact is entailed by the generated text and check hallucinations by inferring the generated text from the input.
LM-based
Filippova (2020); Tian et al. (2020) are based on the intuition that when an unconditional LM, only trained on the targets, gets a smaller loss than a conditional $LM_{x}$, trained on both sources and targets, the token is predicted unfaithfully. Thus, they calculate the ratio of hallucinated tokens to the total target length to measure the hallucination level.
10.3. Hallucination Mitigation in Data-to-Text Generation
Data-Related Methods
Several clean and faithful corpora are collected to tackle the challenges from data infidelity. TOTTO (Parikh et al., 2020) is an open-domain faithful table-to-text dataset, where each sample includes a Wikipedia table with several highlighted cells and a description. To ensure that targets exclude hallucinations, the annotators revise existing Wikipedia candidate sentences and clear the parts unsupported by the table. Moreover, RotoWire-FG (Fact-Grounding) (Wang, 2019) is a purified and enlarged and enriched version of RotoWire (Wiseman et al., 2017) generating NBA game summaries from score tables. Annotators trim the hallucination part in target texts and extract the mapped table records as content plans to better align input tables and output summaries.
For data processing, OpAtt (Nie et al., 2018) designs a gating mechanism and a quantization module for the symbolic operation to augment the record table with pre-calculated results. Nie et al. (2019) utilize a language understanding module to improve the equivalence between the input MR and the reference utterance in the dataset. They train an NLU model with an iterative relabeling procedure: First, they train the model on original data; parse the MR by model inference; train the model on new paired data with high confidence; and then repeat the above processes. Liu et al. (2021b) select training instances based on faithfulness ranking. Finer-grained than the above instance-level method, Rebuffel et al. (2022) label tokens according to co-occurrence analysis and sentence structure through dependency parsing in the pre-processing step to explicate the correspondence between the input table and the text. Generally, the data-related methods are appropriate when the training dataset is noisy.
Modeling and Inference Methods
Planning and skeleton generation are common methods to improve the faithfulness to the input in data-to-text tasks. Liu et al. (2021b) propose a two-step generator with a separate text planner augmented by auxiliary entity information. The planner predicts the plausible content plan based on the input data. Then, given the above input data and the content plan, the sequence generator generates the text. Similarly, Plan-then-Generate (Su et al., 2021) also consists of a content planner and a sequence generator. In addition, this work adopts a structure-aware RL training to generate output text following the generated content plan faithfully. Puduppully and Lapata (2021) first induce a macro plan consisting of multiple sequences of entities and events from the input table and its corresponding multi-paragraph long document. The predicted macro plan then serves as the input to an encoder-decoder model for surface realization. SANA (Wang et al., 2021a) is a skeleton-based two-stage model that includes skeleton generation to select key tokens from the source table and edit-based generation to produce texts via iterative insertion and deletion operations. In contrast to the above two-step model using planning or skeleton, AGGGEN (Xu et al., 2021a) is an end-to-end model that jointly learns to plan and generate at the same time. This architecture with a Hidden Markov Model and Transformer encoder-decoder reintroduces explicit sentence planning stages into neural systems by aligning facts in the target text to input representations.
Other modeling methods have also been proposed to mitigate the hallucination problem. Conjecturing that hallucinations can be caused by inattention to the source, Tian et al. (2020) propose a confidence score and a variational Bayes training framework to learn the score from data. Wang et al. (2020b) introduce a table-text optimal-transport matching loss and an embedding similarity loss to encourage faithfulness. The hallucination degree can also be treated as a controllable factor in generating texts. In Filippova (2020), the hallucination degree of each training sample is estimated and converted into a categorical value which is a part of the inputs as a controlled setting. This approach does not require the dismissal of any input or modification of the model structure.
To mitigate hallucinations at the inference step, Rebuffel et al. (2022) propose a Multi-Branch Decoder that leverages word-level alignment labels between the input table and paired text to learn the relevant parts of the training instance. These word-level labels are gained through dependency parsing during the pre-processing step. The branches separately integrate three co-dependent control factors: content, hallucination, and fluency. Uncertainty-aware beam search (UABS) (Xiao and Wang, 2021) is an extension to beam search to reduce hallucination. Considering that the hallucination probability is positively correlated with predictive uncertainty, this work adds a weighted penalty term in the beam search which is able to balance the predictive probability and uncertainty. This approach is task-agnostic and can also be applied to other tasks, such as image captioning.
These various types of methods do not necessarily conflict and can collaborate to solve the hallucination problem in data-to-text generation.
10.4. Future Directions in Data-to-Text Generation
Given the challenges brought by the discrepancy between structure data and natural text, and the low fault tolerance in the Data-to-Text Generation task, there are several potential directions worth exploring in terms of hallucination.
Firstly, numbers contain information about scales and are common and crucial in the Data-to-Text task  (Suadaa et al., 2021; Zhang et al., 2020b). It is frequent to have errors in numbers, which results in hallucinations and infidelity. This is a serious problem for Data-to-Text generation, yet models rarely give special consideration to the numbers found in the table or text (Thawani et al., 2021). The current automatic metrics of hallucinations also do not specifically treat numbers. This indiscriminate treatment contradicts findings in cognitive neuroscience, where numbers are known to be represented differently from lexical words in a different part of the brain (Göbel and Rushworth, 2004). Thus, considering or highlighting numbers when mitigating and assessing hallucinations is worth exploring. This requires the generative model to learn a better numerical presentation and capture scales, which will reduce the hallucinations caused by the misunderstanding of numbers.
Moreover, for the logical data-to-text generation task, rather than surface-level generation, logical inference, calculation, and comparison are required, which is challenging and causes hallucinations more easily. Thus, reasoning (including numerical reasoning), which is usually combined with graph structures (Chen et al., 2020) is another direction to improve the accuracy of entity relationships and alleviate hallucinations.
11. Hallucinations in Neural Machine Translation
Neural Machine Translation (NMT) is the task of generating translation of the source language into the target language via inference, given parallel data samples for training. Compared to statistical machine translation (SMT) the output of NMT is usually quite fluent and of human-level quality, which creates the danger of misinforming users when there are hallucinations (Martindale et al., 2019).
11.1. Hallucinations Definition and Categories in NMT
The problem of hallucination was identified with the deployment of the first NMT models. Early work comparing SMT and NMT systems (Koehn and Knowles, 2017b), without explicitly using the term “hallucination”, mentioned that NMT models tend to “sacrifice adequacy for the sake of fluency” especially when evaluated with out-of-domain test sets. Following further development of NMT, most of the relevant research papers agree that translated text is considered a hallucination when it is completely disconnected from the source  (Lee et al., 2019; Müller et al., 2020). The categorization of hallucination in NMT is unlike that in any other NLG tasks, and uses various terms that are often overlapping. In order to maintain consistency with other NLG tasks, in this section we use the intrinsic and extrinsic hallucination categories applied to the NMT task by (Zhou et al., 2021b). After a formal definition, we will describe other identified types of hallucinations and hallucination categories mentioned in the relevant literature.
Intrinsic and Extrinsic Hallucinations
Following the idea that hallucinations are outputs that are disconnected from the source, (Zhou et al., 2021b) suggest categorizing the hallucinatory content based on the way the output is disconnected:
  • Intrinsic Hallucination: translations contain incorrect information compared to information present in the source. In Table 3, the example of such hallucination is “Jerry doesn’t go”, since the original name in the source is “Mike” and the verb “to go” is not negated.
  • Extrinsic Hallucination: translations produce additional content without any regard to the source. In Table 3, “happily” and “with his friend” are the two examples of the hallucinatory content since they are added without any apparent connection to the input.
Other Categories and Types of Hallucinations
Raunak et al. (2021) propose an alternative categorization of hallucinations. They divide hallucinations into hallucinations under perturbations and natural hallucinations. Hallucinations under perturbation are those that can be observed if a model tested on the perturbed and unperturbed test set returns drastically different content. Their work on hallucinations under perturbation strictly follows the algorithm proposed by Lee et al. (2019); see Section 11.2.2 on the entropy measure. The second category, natural hallucinations, are created with a connection to the noise in the dataset and can be further divided into detached and oscillatory, where detached hallucinations mean that a target translation is semantically disconnected from a source input, and oscillatory hallucinations mean those that are decoupled from the source by manifesting a repeating n-gram. Tu et al. (2016) and Kong et al. (2019) analyze this phenomenon under the name “over-translation”, that is, a repetitive appearance of words that were not in the source text. Conversely, under-translation is skipping the words that need to be translated (Tu et al., 2016). Finally, abrupt jumps to the end of the sequence and outputs that remain mostly in the source language are also examples of hallucinatory content (Lee et al., 2019).
Category Source Correct Translation Hallucinatory Translation
Intrinsic 迈克周四去书店。
Mike goes to the bookstore on Thursday.
Jerry doesn’t go to the bookstore on Thursday.
Extrinsic 迈克周四去书店。
Mike goes to the bookstore on Thursday.
Mike happily goes to the bookstore on Thursday with his friend.
Detached
Das kann man nur feststellen, wenn die kontrollen mit einer großen intensität durchgeführt werden.
This can only be detected if controls undertaken are more rigorous.
Blood alone moves the wheel of history, i say to you and you will understand, it is a privilege to fight.
Oscillatory
1995 das produktionsvolumen von 30 millionen pizzen wird erreicht.
1995 the production
reached 30 million pizzas.
The US, for example, has been in the past two decades, but has been in the same position as the US, and has been in the United States.
Table 3. Categories and examples of hallucinations in MT by Zhou et al. (2021b) and Raunak et al. (2021)
11.2. Hallucination Metrics in NMT
The definition of hallucinations in machine translation (MT) tends to be qualitative and subjective, and thus researchers often identify hallucinated content manually. Most detrimentally, the appearance of hallucinations is found not to affect the BLEU score of the translated text (Tian et al., 2020; Zhou et al., 2021b). There are, nevertheless, several notable efforts to automatize and quantify the search for hallucinations using statistical methods.
11.2.1. Statistical Metrics
Martindale et al. (2019) propose identifying sentence adequacy using the bag-of-vectors sentence similarity (BVSS) metric. This metric indicates that the information is lost because the reference contains more information than the MT output, or the MT output contains more information than the reference.
11.2.2. Model-Based Metrics
Auxiliary Decoder
“Faithfulness” refers to the amount of source meaning that is faithfully expressed in the translation, and it is used interchangeably with the term “adequacy” (Tu et al., 2017; Feng et al., 2020). Feng et al. (2020) propose adding another “evaluation decoder” apart from the standard translation decoder. In their work, faithfulness is based on word-by-word translation probabilities, and is calculated in the evaluation module along with translation fluency. The loss returned by the evaluation module helps to adjust the probability returned by the translation module.
Entropy Measure
In scenarios where the ground truth of a translation is not available, an entropy measure of the average attention distribution can be used to detect hallucinations. Tu et al. (2016) and Garg et al. (2019) show that hallucinations are visible in attention matrices. When the model outputs correct translation, the attention mechanism attends to the entire input sequence throughout decoding. However, it tends to concentrate on one point when the model outputs hallucinatory content. The entropy is calculated on the average attention weights when the model does or does not produce hallucinations during testing. For comparison, a clean test set is used along with the purposefully perturbed one, which is created to incite hallucinations (test sets featuring multiple repetitions). The mean entropy returned by hallucinatory models diverges from the mean of the models that do not produce hallucinations spontaneously (Lee et al., 2019).
Token Level Hallucination Detection
Zhou et al. (2021b) propose a method for detecting hallucinated tokens within a sentence, making the search more fine-grained. They use a synthetic dataset that is created by adding noise to the source data, more specifically it is generated by a language model with certain tokens of correct translations masked. Tokens in synthetic data are labeled as hallucinated (1) or not (0). Then the authors compute the hallucination prediction loss between binary labels and the tokens from the hallucinated sentence. This work further employs a word alignment-based method and overlap-based method as baselines for hallucination.
Similarity-based Mathods
Zhou et al. (2021b) use an unsupervised model that extracts alignments from similarity matrices of word embeddings (Sabet et al., 2020), and then predicts the target token as hallucinated if it is not aligned to the source. Parthasarathi et al. (2021) propose calculating faithfulness by computing similarity scores between perturbed source sentence and target sentence after applying the same perturbation.
Overlap-based Mathods
Zhou et al. (2021b) predict that the target token is hallucinated if it does not appear in the source. Since the target and source are two different languages, the authors use the density matching method for bilingual synonyms from Zhou et al. (2019). Kong et al. (2019) suggest the Coverage Difference Ratio (CDR) as the metric to evaluate adequacy, which is especially successful in finding cases of under-translation. It is estimated by comparing source words covered by generated translation with human translations.
The overlap-based methods for detecting hallucinations are heuristics based on the assumption that all translated words should appear in the source. However, this is not always the case, e.g., when paraphrasing or using synonyms. Using word embeddings as similarity-based methods helps avoid such simplifications and allows more diverse, synonymous translations.
Approximate Natural Hallucination Detection
Raunak et al. (2021) propose Approximate Natural Hallucination (ANH) detection based on the fact that hallucinations often occur as oscillations (repeating n-grams) and the lower unique bigram count indicates a higher appearance of oscillatory hallucinations. Furthermore, the ANH detection method searches for repeated targets in the translation output. Their method finds translation above a certain n-gram threshold and searches for repeated targets in the output translation, following the assumption that if hallucinations are often incited by aligning unique sources to the same target, then repeating targets will also appear during the inference (Tu et al., 2016).
11.3. Hallucination Mitigation Methods in NMT
Hallucinations in MT are hard to discover for a person who is not fluent in the target language, and thus they can lead to many possible errors, or even dangers. Out of all the natural language generation tasks, NMT engines such as Google in the English-speaking internet and Baidu in the Sinosphere are probably the most widely accessible to netizens. Consequently, there is a big interest in improving NMT´s performance, also by mitigating hallucinations. This subsection compiles methods of mitigating hallucinations in NMT.
11.3.1. Data-Related
Data augmentation appears to be one of the most common methods for removing hallucination. Lee et al. (2019) and Raunak et al. (2021) suggest addition of perturbed sentences. Furthermore, perturbation, where the insertions of most common tokens are placed at the beginning of the sentence, seems to be the most successful in hallucination mitigation. A disadvantage of this method is the need to understand different types of hallucinations produced by the model in order to apply a correct augmentation method. Corpus filtering is a method of mitigating hallucinations caused by the noise in the dataset by removing the repetitive and mismatching source and target sequences (Raunak et al., 2021). Junczys-Dowmunt (2018) implements a cross-entropy data filtering method for bilingual data, which uses cross-entropy scores calculated for noisy pairs according to two translation models trained on the clean data. The scores that suggest dissagreament between sentence pairs from two models are subsequently penalized.
While (Raunak et al., 2021; Lee et al., 2019) and (Junczys-Dowmunt, 2018) define noise as mismatched source and target sentences,  (Briakou and Carpuat, 2021) analyzes the influence of fine-grained semantic divergences on NMT outputs. The authors consequently propose a mitigation method for fine-grained divergences based on semantic factors. The tags are applied to each source and target sentence to inform about the position of divergent tokens. Factorizing divergence not only helps to mitigate hallucinations, but improves the overall performance of the NMT. This shows that tagging small semantic divergences can provide useful information for the network during training.
11.3.2. Modeling and Inference
Overexposure bias is a common problem in NMT, amplified by the teacher-forcing technique used in sequence-to-sequence models. The models are trained on the ground truth, but during inference, they attend to the past predictions, which can be incorrect (Ranzato et al., 2016; Kong et al., 2019). To mitigate this problem, Wang and Sennrich (2020) propose substituting MLE as a training objective with minimum risk training (MRT) (Och, 2003). Scheduled sampling is a classic method of mitigating overexposure bias first proposed by (Bengio et al., 2015). Based on that method, (Goyal et al., 2017) create a differentiable approximation to greedy decoding that shows a good performance in the NMT task. (Xu et al., 2019) propose further improvement of the scheduled sampling algorithm for NMT by optimizing the probability of source and target word alignments. This improvement helps to address the issue flexibility in word order between a source and target language when performing scheduled sampling.
Zhou et al. (2021b) propose a method of improving self-training of NMT based on hallucination detection. They create hallucination labels (see Section 11.2.2), and then discard losses of tokens predicted as hallucinations, which is known as token loss truncation. This is similar to the method proposed by Kang and Hashimoto (2020a), the latter for full sentences in the summarization task. Furthermore, instead of adjusting losses, Zhou et al. (2021b) mask the hidden states of the discarded losses in the decoder in a procedure called decoder HS masking. Experimental results show both a translation quality improvement in terms of BLEU and also a large reduction in hallucination. The token loss truncation method shows good results in the low-resource languages scenario.
Another method to mitigate the impact of noisy datasets is tilted empirical risk minimization (TERM), a training objective proposed by Li et al. (2020a). (Lee et al., 2019) mentions that techniques such as dropout, L2E regularization, and clipping tend to decrease the number of hallucinations. Lastly, several authors propose methods of improving phrase alignment that are helpful both in increasing translation accuracy and identifying content that did not appear in the source translation (Weng et al., 2020; Zhang et al., 2021b; Garg et al., 2019).
11.4. Future Directions in NMT
The future work on hallucinations in NMT is to define hallucinations in a quantifiable manner; i.e., to specify a cut-off value between translation error and hallucinated content using a particular metric. Martindale et al. (2019) propose a threshold between fluency and adequacy which is the closest to this ideal. They, however, do not concentrate on hallucinated content as such, and thus fluent but inadequate sentences may not always indicate hallucinations but also other types of translation errors. Balakrishnan et al. (2019) mention constrained decoding as a method to mitigate hallucinations in dialogue systems, but it could also be applied in NMT. (Hokamp and Liu, 2017; Post and Vilar, 2018; Song et al., 2019; Dinu et al., 2019; Susanto et al., 2020; Xu and Carpuat, 2021) and (Xu and Carpuat, 2020) use constrained decoding to incorporate specific terminology into MT, but the above methods can be repurposed to mitigate hallucinations.
Another direction for future work on hallucinations is improving existing methods of searching for hallucinatory content, such as the algorithms proposed by Lee et al. (2019); Feng et al. (2020) and Raunak et al. (2021), that are computationally expensive (Raunak et al., 2021) or require the creation of an additional perturbed test-set (Lee et al., 2019). Similarly, for mitigation of lack of faithfulness and fluency, the method proposed by Feng et al. (2020) requires the creation of a one-to-many architecture (one encoder and two decoders), which is also computationally expensive. Future directions would therefore include simplification of existing hallucination evaluation methods, applying them to different architectures like CNNs and transformers, and possibly conducting research on finding simpler hallucination search methods.
12. Hallucination in Vision-Language Generation
With the vast advancement of the Transformer architecture (Vaswani et al., 2017; Dosovitskiy et al., 2021) in both CV and NLP, there is a trend to pre-train large-scale unified vision-language (VL) models (Wang et al., 2022b; Dai et al., 2022a; Li et al., 2022; Wang et al., 2022c; Alayrac et al., 2022; Wang et al., 2022a) to perform vision grounded text generation tasks, such as image captioning and visual question answering (VQA). Generally, there are two common schemas for vision-language pre-training: 1) pre-train from scratch with a massive amount of image-text pairs as well as optionally a large text-only corpus; or 2) initialize model parameters from a large pre-trained LM and then adapt it to the VL domain with adequate image-text pairs. Either way, the learned vision and language representations are aligned in the same multimodal space and the resulting model can be seen as a LM that understands visual information. Therefore, the hallucination problem is also observed in VL models due to similar reasons as found in NLG.
In the VL domain, the research on hallucination is still in its very early stage and how to measure and mitigate hallucination is an open question. In this section, we first review the hallucination in image captioning as it is the only VL task that has corresponding previous research works. Then, we introduce hallucination phenomena found in other VL tasks. Finally, we discuss potential future research directions on this problem.
12.1. Object Hallucination in Image Captioning
Definition.
Object hallucination is defined as models generating captions that contain non-existent or inaccurate objects from the input image. Following tasks in NLG, we also categorize object hallucination into intrinsic and extrinsic ones:
Figure 2. Examples of intrinsic and extrinsic object hallucination in image captioning.
  • Intrinsic Object Hallucination: captions contain incorrect or definitely non-existent objects given the input image. For example in Figure 2, there is no “mirror” or “football” on top of the chest in the given image.
  • Extrinsic Object Hallucination: captions contain objects cannot be verified their existence from the input image. For example in Figure 2, we cannot verify whether there are “letters” in the drawer or a “fan” on the roof.
Metrics
To automatically measure object hallucination, Rohrbach et al. (2018) propose the CHAIR (Caption Hallucination Assessment with Image Relevance) metric, which calculates what proportion of object words generated are actually in the image according to the ground truth captions. Specifically, there are two variants of it, which are CHAIRi and CHAIRs defined as follows,
[…]
CHAIRi measures per-instance object hallucination, i.e. what fraction of object instances in each generated caption are hallucinated. CHAIRs measures per-sentence object hallucination, i.e. what fraction of generated captions include at least one hallucinated object. For example, to calculate CHAIR scores for the MSCOCO dataset (Lin et al., 2014), Rohrbach et al. (2018) apply the 80-object list used in the MSCOCO segmentation challenge Lu et al. (2018) and find exact matches of object words or phrases in captions.
Mitigation
As a research problem that is in its early stage, there are currently a limited number of approaches proposed to mitigate object hallucination in image captioning. Biten et al. (2022) hypothesize that the main cause of object hallucination is the systematic co-occurence of particular object categories in input images. They propose three simple yet effective ways of data augmentation to make the co-occurence statistics matrix more uniform to mitigate object hallucination. Results show that their introduced method can reduce object hallucination without changing model architectures. From another perspective, Xiao and Wang (2021) propose an uncertainty-aware beam search method for decoding and exhibit that reducing uncertainty can lead to less hallucination. Specifically, a weighted penalty term is added to the beam search objective to balance between log probability and predictive uncertainty of the selected word candidates. More recently, Dai et al. (2022b) analyze object hallucination in VL pre-training and propose a novel pre-training objective named object masked language modeling to alleviate this problem.
12.2. Hallucination in Other VL Tasks
In addition to image captioning, hallucination has also been observed in other VL tasks and raised as an open research question. For example, in open-ended visual question answering, Figure 3 (left and right) shows that the model could generate seem likely answers when we only see the text, however wrong when given the image. Moreover, Figure 3 (middle) indicates that hallucination can also be triggered by adversarially prompting an unanswerable question. The model will imagine an unsupported answer that commonly matches the given visual scene.
Figure 3. Examples of hallucination in visual question answering (taken from (Alayrac et al., 2022)). The bold text is the output generated by the model and the part before it is the input prompt.
12.3. Future Directions in VL
For future research on the hallucination problem in VL, we summarize three promising directions. Firstly, hallucination detection and mitigation in VL is still in the early stage. There is a lack of empirical and theoretical analyses in many tasks, such as visual storytelling, visual commonsense reasoning, video captioning, etc. Secondly, more effective evaluation metrics are needed. For example, although CHAIR can automatically evaluate the degree of object hallucination in image captioning, it requires a pre-defined list of object categories, which does not generalize well. Furthermore, currently there is no automatic metric for the hallucination types discussed in Section 12.2. Therefore, we cannot perform quantitative evaluations for them. Thirdly, we believe how to perform controlled generation (Rebuffel et al., 2022; Dai et al., 2022b) with visual grounding is a promising direction to mitigate hallucination in VL.
13. Hallucination in Large Language Models
Scaling up model and data size for language models has shown great empirical success. The resulting Large Language Models (LLMs) not only have achieved significant performance improvements upon previous PLMs across a diverse array of NLP tasks, but have also demonstrated many emerging abilities and strong steerability after instruction tuning and Reinforcement Learning from Human Feedback (RLHF) (Zhao et al., 2023; Wei et al., 2022). However, they still exhibit the hallucination problem. Even worse, since LLMs generate highly fluent and convincing responses, their hallucinations become more difficult to identify, and more likely to have harmful consequences. The launch of ChatGPT as an LLM with a conversational user interface led to the popularity of LLMs and their wide range of real-world applications. The research on the identification and mitigation of hallucination in LLMs has intensified. LLMs are usually open-ended general-purpose systems, which differ much from task-specific models, and their architectural designs, data coverage, training methodologies, and model behaviors are also different from PLMs mentioned in previous sections. Therefore, in the following sections, we specifically discuss the hallucination problem in LLMs, by covering more recent works, introducing novel strategies for measuring and mitigating hallucination, as well as listing unsolved questions and future directions.  [^8] This section was updated in Jan 2024.
13.1. Hallucination Definition in LLMs
In the realm of LLMs, hallucination takes on a broader definition, reflecting the vastness of their training data, the breadth of their knowledge base, and their multitasking capability. Hallucination in LLMs not only signifies deviations from the source input but also extends to deviations from world knowledge. In this context, the “fact” discussed in Section 2.3 includes both the input source and the world knowledge. The hallucination degree reflects and encapsulates the model’s capacity to accurately and faithfully comprehend and represent the world. This includes a grasp of factual information, an understanding of linguistic subtleties, and an ability to navigate the intricate labyrinth of human knowledge and experience. Therefore, LLM hallucination is more oriented toward the extrinsic type involving unfaithful or nonsensical facts. Researchers explore the model’s ability to maintain the integrity of its parameterized knowledge, contextual understanding, recall relevant knowledge, and reasoning abilities (Ji et al., 2023c). This definition of hallucination in LLMs underscores the challenges and opportunities in developing these models. It highlights the need for robust mechanisms to prevent hallucination, to ensure the accuracy and reliability of the models, and to harness their potential in a responsible and beneficial manner.
13.2. Hallucination Metrics for LLMs
Hallucination metrics for LLMs are designed to either measure the level of hallucination in LLM-generated content, or quantify the hallucination risk of individual LLMs. Most of the related work focuses on measuring extrinsic hallucinations in the open domain. Depending on whether a metric requires golden references to identify hallucinations, we divide existing hallucination metrics for LLMs into reference-dependent and reference-free ones. In reference-dependent metrics, generations of LLM are analyzed based on references sourced from verified knowledge bases such as Wikipedia, news articles, books, and knowledge graphs. For reference-free metrics, hallucinations are identified by token probability-based uncertainty measures or self-consistency checks.
13.2.1. Reference-dependent Metrics
Reference-dependent metrics are often formatted as “hallucination benchmarks” that measure the hallucination level not risks of different LLMs. Most benchmarks utilize multiple choice QA or fill-in-the-blank QA to assess model hallucination, while some efforts further consider the ratio of partially incorrect answers (Sun et al., 2023) or instances where the model admits that it cannot answer (Lin et al., 2022). Some benchmarks adopt the open domain and open-ended QA setting, where exact-match-based evaluations generally fail to recognize semantically equivalent answers (Kamalloo et al., 2023). Free-form responses from LLMs can be analyzed based on atomized decomposition (FactualityPrompt (Lee et al., 2022) and FActScore (Min et al., 2023)), LLM-based evaluation (self-judgment (Ren et al., 2023) or GPT-4-based judgment (Adlakha et al., 2023)), or by measuring the discrepancy of the vanilla response and the response giving the knowledge for reference (Yu et al., 2023a; Huo et al., 2023; Pezeshkpour, 2023).
Human-annotated Hallucination Benchmarks
Collecting human-verified QA samples to test the factual knowledge of LLMs is the most straightforward method. Since the training corpora of LLMs are extremely large, high frequency count knowledge and basic commonsense can be easily memorized. As a result, examining low frequency, long-tailed knowledge becomes a primary focus of these hallucination benchmarks. Testing samples for long-tail knowledge can be selected based on frequency of appearance (e.g., in pretraining corpora (Kandpal et al., 2023) or Wikitext (Du et al., 2023)), popularity measurements based on website traffic data (e.g., PopQA (Mallen et al., 2023), Head-to-Tail (Sun et al., 2023)), recency (RealTimeQA (Kasai et al., 2022)), or from specific domains (e.g., ExpertQA (Malaviya et al., 2023) and Med-HALT (Umapathi et al., 2023)). In addition to long-tail knowledge, another way to build challenging benchmarks is to gather questions that are more likely to lead to hallucinations – a methodology similar to adversarial prompting (Kumar et al., 2023; Zhu et al., 2023) and red teaming (Ganguli et al., 2022). The most representative example is TruthfulQA (Lin et al., 2022), a benchmark containing challenging questions that some humans would answer falsely due to false beliefs or misconceptions (e.g., question “If it’s cold outside, what does that tell us about global warming?” leads to GPT-3 answers “It tells us that global warming is a hoax”). Similarly, the DecodingTrust (Wang et al., 2023a) evaluated the trustworthiness of GPT models in terms of adversarial robustness, OOD knowledge, vulnerability of jail-breaking, misleading instructions, etc. Known-Unknown (Amayuelas et al., 2023) and SeflAware (Yin et al., 2023b) evaluated LLMs’ awareness of uncertainty for question without definitive answer (e.g., “If the Universe started at the Big Bang, what existed before then?”)
Automated Hallucination Benchmarking
Manual question collection and answer annotation are labor intensive, hard to scale, and hard to adapt to new domains. To address this issue, annotation-free benchmarking methods have been proposed, which can be further divided into 1) automated benchmark generation and 2) automated evidence retrieval. For the first type of methods (Muhlgay et al., 2023; Chen et al., 2023c), factual but unstructured corpora (for example, Wikipedia and reliable news articles) are transformed into a unified QA format based on information extraction and answer candidate generation. The second type of methods automatically retrieves evidence related to LLM generations from the Internet, similar to automated fact checking (Thorne and Vlachos, 2018; Guo et al., 2022). They typically adopt a chained pipeline, where search queries are first generated (by an LLM) based on the content that needs to be verified, and then a model determines whether it contains hallucinations according to the retrieved evidence (Gao et al., 2023)). This pipeline can be further enhanced by upgrading the verification model from an NLI model (Gao et al., 2023) to prompted/fine-tuned LLMs (as in AttributionScore (Yue et al., 2023) and (Varshney et al., 2023a)), or by adding additional steps to the pipeline, such as tool interaction (as in FactTool (Chern et al., 2023) and CRITIC (Gou et al., 2023a)), claim decomposition (Self-Checker (Li et al., 2023e)), additional second stage fine-grained retrieval and claim-focused summarization (as in (Chen et al., 2023a)), or chain-of-verification (Dhuliawala et al., 2023a).
13.2.2. Reference-free Metrics
Reference-free metrics aim to assess the hallucination risk of LLMs without relying on external references or ground truth. Although not as reliable as reference-based ones, these metrics are particularly useful when golden references are unavailable, expensive to obtain, or when hallucination identification needs to be performed in real time. Reference-free metrics generally fall into two categories: uncertainty-based and consistency-based hallucination detection. Uncertainty-based methods rely on the token probabilities assigned by the LLM during generation, whereas consistency-based methods evaluate the coherence of multiple completions generated by the LLM. There are also some benchmarks to measure the effectiveness of these reference-free hallucination metrics, which are also termed “hallucination detection benchmarks”, such as HaluEval (Li et al., 2023b) and HaDes (Liu et al., 2022) These benchmarks serve as meta-metrics of hallucination metrics, offering valuable resources for developing more effective methods to identify hallucinations.
Uncertainty-based Hallucination Detection
This type of method assumes that LLMs assign high probabilities to tokens that they are confident of, and low probabilities to uncertain tokens, which usually contain hallucinated information. It has a close relationship with classical uncertainty estimation methods (Abdar et al., 2020; Mena et al., 2022), since the generation of each token can be viewed as a vocabulary size classification problem. The difference here is that autoregressive text generation of LLM is a chained classification process, therefore individual token (subword) probabilities need to be aggregated so that they can reflect word-/sentence-/passage-level uncertainty. Such aggregation can be done by average-/max-/min-pooling on token probabilities, calculating the normalized product of the token probabilities, or calculating the maximum/averaged entropy (Manakul et al., 2023; Varshney et al., 2023b; Huang et al., 2023; Kadavath2022Language). These methods can be further extended by adding prompts like “Generate factually consistent summary for the following text: ¡source-text¿ ¡generated-text¿” (Fu et al., 2023) for fine-grained control (factuality evaluation on summarization), or “¡question¿ ¡generated-answer¿ The answer is true.” (Kadavath2022Language) for self-evaluation of QA samples.
Consistency-based Hallucination Detection
is the more reliable and common approach. In practical applications, many LLMs only offer an API while internal logits remain inaccessible (i.e., the “black-box setting”), thus making uncertainty-based methods inapplicable. Efforts have been made to create surrogate approximations of token probability. Specifically, LLMs first generate multiple completions with stochastic sampling given a fixed context, and then the consistency of these completions is used to reflect the uncertainty. Consistency can be measured by the BLUE-based variation ratio (Huang et al., 2023), BERTScore (Manakul et al., 2023; Raj et al., 2022), n-gram approximation (Manakul et al., 2023), NER-based overlap ratio (Raj et al., 2022), NLI model (Raj et al., 2022; Elaraby et al., 2023), or LLM-based judgment (Mündler et al., 2023).
13.3. Hallucination Mitigation in LLMs
13.3.1. Data-Related Methods
Data for Pretraining
Recent studies (Zhou et al., 2023) showed that the knowledge within LLMs is almost entirely acquired during pretraining. Accordingly, a strong emphasis should be placed on ensuring the quality of the pretraining data. This can be performed through the collection of pretraining data from credible sources (Touvron et al., 2023a) and the minimization of defective and noisy data such as those that are unreliable or unverifiable (Zhang et al., 2023a). A representative example of such data-driven methods can be found in the development of Llama-2 (Touvron et al., 2023b), where the most factual sources within the pretraining data were up-sampled in order to reduce hallucinations. Similarly, Falcon LLM (Penedo et al., 2023) also demonstrated that their data refinement methods including rigorously filtering and deduplicating can significantly boost LLM performance. Due to the extensive pretraining datasets, some approaches have been developed to overcome the impracticalities that arise from manually selecting them. For instance, for the pretraining of GPT-3 (Brown et al., 2020), a classifier-based automatic filtering method was applied to remove low-quality documents. However, as pretraining data is being continuously scaled up, it is becoming increasingly costly for the research community to even access the entire pretraining dataset. As an alternative, small LLMs like phi-1.5 (Li et al., 2023a), which are trained on strictly controlled “textbooks-style” small-scale corpus, provide a valuable opportunity for doing rigors ablations to investigate the relationship between pretraining data and model hallucination.
Data for Instruction-tuning
During the fine-tuning stage, data refinement can also be used to dampen hallucinations. As the size of fine-tuning data is significantly smaller, manual approaches can also be practically used (Zhang et al., 2023a). For example, the authors of LIMA collected a diverse dataset of 1,000 carefully curated prompts and responses which are aligned with each other (Zhou et al., 2023). Conversely, Chen et al. (2023b) point out that widely used fine-tuning datasets e.g. Alpaca’s 52k data contain low-quality instances with incorrect or irrelevant responses which are detrimental to fine-tuning, used an automated approach leveraging strong LLMs to identify and discard such low-quality data. Alongside, the authors of InstructMining (Cao et al., 2023), proposed linear rules for selecting high-quality instruction fine-tuning data and avoided the need for human or machine annotations.
Data for Reward Model Training
Training reliable reward models such that desirable and undesirable outputs are distinguished is an effective method in mitigating hallucination. These models can subsequently be incorporated into reinforcement learning pipelines. Since the effectiveness of the system is constrained by the performance of the reward model itself, emphasis has been placed on determining the most effective training methods (Lightman et al., 2023). Thus, OpenAI released a dataset, namely, PRM800K, which resulted in a SOTA reward model (Lightman et al., 2023). In the development of GPT-4 (OpenAI, 2023), two different approaches were implemented. Firstly, in tackling open-domain hallucinations, OpenAI collected real-world ChatGPT data that were flagged as non-factual by users and used it together with additional labelled comparison data to train their reward models. Secondly, for closed-domain hallucinations, OpenAI utilised GPT-4 itself to generate and subsequently utilise synthetic data for reward model training.
Information Augmentation
Hallucinations can be mitigated via the use of external knowledge, or tools. Retrieval Augmented Generation (RAG) is a framework that utilizes data to improve the quality of LLM performances through grounding the model with an external knowledge base (Martineau, 2023; Zhang et al., 2023b) which is a popular method used to mitigate hallucination (Es et al., 2023). However, although being a common approach, Zhang et al. (2023b) argue that despite improving end-task performance, RAG does not consistently mitigate the issue of hallucinations due to misalignment between user questions and stored knowledge. The LLM may, as a result, draw invalid correlations or utilize biased model knowledge to override retrieved knowledge and subsequently hallucinate. To enhance this alignment, automatic question refinement such as MixAlign (Zhang et al., 2023b) achieves this question-knowledge alignment through the use of a language model. Furthermore, it can also achieve further enhancement, if necessary, through human user clarifications. More specifically, the framework employs the LLM to map the conditions and constraints of the user question to corresponding ones in the knowledge base and in cases where this mapping process yields uncertainties or the evidence remains elusive, MixAlign generates a question seeking further clarification from the user. This approach has been shown to not only decrease model hallucinations but also improve the quality of generated responses.
Examples of leveraging external knowledge for answer refinement REFEED (Yu et al., 2023c). REFEED is an approach that first generates initial outputs, uses a retrieval model to acquire information from large document collections, and lastly refines the initial output by incorporating this retrieved information, and LLM-Augmenter, which is a plug-and-play method that uses automatic feedback. CRITIC (Gou et al., 2023b) is an example of a framework that uses external tools such as search engines and code interpreters to verify the initial output and successively amend it based on the critiques of the verification.
13.3.2. Modelling and Inference Methods
Reinforcement Learning with Human Feedback and Safety Fine-Tuning
Reinforcement learning with human feedback (RLHF) and safety fine-tuning are one of the most dominant methods utilized to mitigate hallucination. One notable example is the resultant model of GPT-3 having undergone fine-tuning and further fine tuning using supervised learning and RLHF respectively, coined as InstructGPT, which roughly halved the hallucination rate from 41% to 21% in closed domain tasks such as summarization and closed domain QA (Ouyang et al., 2022). In another well-known study, the authors of Llama-2-Chat (Touvron et al., 2023b) observed a compelling finding, which is that after various RLHF iterations, despite the rise in temperature of the LLM, for prompts based on factual information, the model learns to consistently provide the same factual response whilst retaining its capability in generating diversity when fed creativity-based prompts. Furthermore, the authors utilize RLHF with the aim to improve the helpfulness (defined as how well the LLM responses fulfill users’ requests and provide requested information) and safety of the LLM (Touvron et al., 2023b), which can decrease hallucination. Moreover, supervised safety fine-tuning was also carried out to improve model safety. It should be noted that the criterion for safety in this study not only involves whether harmful responses were produced, but also unhelpful ones (Touvron et al., 2023b). As part of RLHF, the reward models were initialized from pretrained chat model checkpoints to ensure that both the chat model and reward model benefit from knowledge acquired form pretraining thereby avoiding an information mismatch which could give rise to hallucinations (Touvron et al., 2023b).
Decoding Methods
The modular architecture of LLMs can also be leveraged to combat hallucination. For example, researchers from MIT and Microsoft propose an approach, namely, Decoding by Contrasting Layers (DoLa), which emphasises knowledge from higher layers of an LLM whilst downplaying lower or intermediate layer knowledge (Chuang et al., 2023). Since factual knowledge is progressively injected by higher layers, the probability of the correct next word increases further up the architecture. Therefore, by exploiting the modularity of knowledge encoding in LLMs, the authors propose to amplify factual knowledge of LLMs via obtaining the probability of the next word from the difference between logits produced by a higher and lower layer. By making this contrast, the prominence of the correct next word’s probability is increased. Results from the study demonstrate that DoLa consistently improves truthfulness which consequently reduces hallucination.
Similarly, Context-Aware Decoding (CAD) (Shi et al., 2023), also exploits logits and uses a contrastive approach to mitigate hallucination. However, instead of contrasting between layers, CAD amplifies the difference between the output probabilities when a model is used with context and without. This approach effectively suppresses prior knowledge when contextual information is provided.
Model Editing
Subtracting the parameters of a “negative expert” fine-tuned on negative samples from a pre-trained LLM can also serve to reduce hallucination (Daheim et al., 2023). However, given that some parameters are more responsible than others in causing hallucination, Daheim et al. (Daheim et al., 2023) proposed an improved method coined as Elastic Weight Removal (EWR), which uses an approximation of the Fisher Information matrix, to weigh their individual performance, and computes a vector of differences between pre-trained and hallucination (anti-)expert models (trained with synthetic data) based on the weighting. Similarly, parameter interpolations with an abstractiveness expert model were also performed to prevent an increase in extractives, which is a potential side-effect of discouraging hallucinations. This steers the model towards being able to re-phrase ground-truth knowledge and drawing new valid conclusions rather than directly copying sections from documents with a limited understanding of their meaning.
Based on the notion that models “know” more than they “say”, the researchers at Harvard (Li et al., 2023d) studied the difference between generation accuracy determined by a model’s output and the probe accuracy (i.e. sentence classification using a model’s intermediate activations as input to a classifier). The difference observed was 40%, which shows a major gap between the information present at intermediate layers and the output. Thus, to narrow this gap, the authors propose Inference-Time intervention, a computationally inexpensive and data-efficient technique that requires only a few hundred samples. The approach involves first using supervised learning to identify attention heads with high linear probing accuracy for truthfulness, in other words, identifying latent vectors related to factual outputs using supervised learning. And secondly, during LLM inference, shift model activations along these truth-correlated directions. The same intervention is repeated autoregressively until the entire answer is generated. ITI was shown to improve truthfulness from 32.5% to 65.1% on Alpaca, an instruction fine-tuned Llama, which suggests the effectiveness of this approach in mitigating hallucinations.
Chain-of-Thought and Variants
Chain-of-Thought can be leveraged to reduce hallucinations. For example, simply adding the prompt, “Let’s think step by step”, encourages LLMs to reason prior to arriving to an answer, which improves model performance (Kojima et al., 2023). This is further evidenced in a related study where error rates were significantly reduced through using the same prompting strategy (Zhang et al., 2023c).
At inference, reasoning-based methods such as Chain-of-Verification (COVE) (Dhuliawala et al., 2023b) and Chain of Natural Language Inference (CoNLI) (Lei et al., 2023) have been developed to reduce hallucination. The COVE method consists of four steps where an LLM first drafts an initial response, plans verification questions to fact check this, answers the questions independently to ensure the answers are not biased by other responses, and finally generates its final verified response. The framework, CoNLI, involves three steps, that is, to firstly select sentences (considered in this case as a set of hypotheses), secondly detect hallucinations with sentence and entity level detectors through prompting the LLM to solve a sequence of inference problems, and lastly to use a mitigation agent which uses detection reasonings as mitigation instructions to produce a refined response.
Self-reflection and self-collaboration is another approach that can be used to combat hallucination (Ji et al., 2023c; Shinn et al., 2023; Wang et al., 2023b). Ji et al. (2023c) is an iterative feedback loop process that involves generating relevant background knowledge for a given question followed by a factuality evaluation. When discrepancies are detected, the model is made to self-correct, which leverages its inherent capacity to reflect and thereby refine the knowledge. This iteration is continued until a satisfactory level of factuality is attained. Furthermore, during the answering stage, a similar generation-score-refine strategy is used along with an entailment evaluation to determine whether the answer properly addresses the question. If the answer fails to do so, the process returns back to the initial stage and repeats the cycle. This dynamic interaction between the system and its knowledge enhances the model to provide accurate, reliable, as well as factually grounded responses. A Solo Performance Prompting (SPP) (Wang et al., 2023b) is proposed in light of LLMs’ tendency to struggle with tasks that require intensive domain knowledge and complex reasoning. Through prompting the LLM to engage in multi-turn self-collaboration with multiple personas, SPP transforms a sole LLM into an intelligent agent that works with multiple minds such that their individual strengths and knowledge are combined in order to boost overall problem-solving and performance in such tasks.
Ensemble
Ensembling multiple models to mitigate hallucination is another way. The “Society of Minds” approach can be used as an alternative to methods that are data-driven, modeling-driven, or at inference. One example of such (multiagentdebate2023), where multiple LLMs individually produce a response and then jointly debate each other on the responses and reasoning, such that a single common answer is reached, has been demonstrated to improve the factuality of generated content and reduce hallucination. This multi-agent interaction more specifically involves multiple instances of an LLM to first generate individual candidate answers to a given query. Subsequently, each individual model instance reads and critiques the responses of all other models and uses this content to refine its own answer; a step which is reiterated several times. This entire process steers the models towards constructing answers that are consistent with their own internal criticism and also sensible provided the response from other agents.
Post-processing
PURR (Petite Unsupervised Research and Revision) (Chen et al., 2023e), requires LLMs to introduce corruptions/noises into data and subsequently fine-tune an editor to denoise these corruptions. Specifically, PURR, a fusion-in-decoder T5 model, is trained to denoise corruptions when simultaneously given relevant evidence. Given an ungrounded output generated by LLMs, PURR generates questions to search for relevant evidence, which is subsequently used to produce an edit for the output.
13.4. Future Directions of Hallucination Mitigation in LLMs
13.4.1. Hallucination in Large Multimodal Models
Large multimodal models (LMMs) such as GPT-4V can generate rich and detailed response for visual inputs, and therefore being inherently riskier compared to previous VLMs for single-sentence captioning and short-answer VQA. Furthermore, images convey different levels of information encompassing object existence, attributes (color, shape, etc), spatial relationships, and sometimes high-level emotional responses they elicit (peaceful, beautiful, etc.), which makes mitigating hallucination in LMM even more challenging. The exploration of this area is still in its early stage, with a primary focus on object existence hallucination (Li et al., 2023c). However, it is imperative for LMMs to analyze images well beyond that surface level, otherwise the use of LMM would be redundant and a simple object detector would suffice. The most crucial factor in this problem is the availability and quality of data and supervision. Language-supervised representations (e.g., CLIP) have demonstrated inherent shortcomings in recognizing certain types of visual patterns, such as object orientation, quantity, and viewpoint (Tong et al., 2024). Even with perfect representation, imperfect vision-language alignment can result in a significant gap between the visual recognition abilities of the LMM and its visual backbone (Zhai et al., 2023). The cost of annotation during the mitigation of multimodal hallucination should also be considered. While there are works on the visual extension of faithfulness-oriented RLHF (Yu et al., 2023b), these essentially involve labeling more data, which is expensive and difficult to scale. The versatility of the visual domains also makes it challenging to guarantee robustness and generalization. Large-scale self-supervision may be the path forward, as supervised learning has proven to be insufficient for both the CV and NLP in the past decade.
13.4.2. Hallucination in Long-tail and Low-resource Domains
Hallucinations in LLMs in long-tail and low-resource domains still remains a significant challenge. This has become a new research focus – for example, many recent hallucination benchmarks are proposed to measure LLM hallucination in various long-tail and low-resource domains (see Section 13.2.1). However, our scope should not be confined to only narrow professional areas (e.g., medical, legal, finance), but should also consider broader scenarios which are also long-tailed and low-resource. For example, hallucination in low-resource languages is a less explored area, as hallucination mitigation methods applied to rich-resource languages do not necessarily generalize to low-resource ones. Similarly, the temporal axis also impacts the frequency of world knowledge appearance, where more recent data could also be considered as long-tailed and low-resource. Furthermore, beyond textual data, multi-modal scenarios are inherently low-resource for text-only pretrained LLMs. In light of the above, instead of developing domain-specific methods for hallucination mitigation, we need to design unified framework and generalizable strategies for addressing hallucination in various imbalanced scenarios.
13.4.3. Estimating the Knowledge Boundary and Expressing the Uncertainty
Despite the vast size and continuous scaling of pretraining datasets for LLMs, it is impossible for them to encompass all the world’s knowledge. Consequently, teaching models to express their inability to answer specific questions and honestly admit“I don’t know” is crucial, and it requires LLMs to accurately model their knowledge boundaries. Notably, this task lies in a different dimension from other objectives that we expect LLMs to achieve (e.g., improving QA task accuracy) since it is not explicitly included in the pretraining corpora, making it more challenging. Furthermore, the diverse sources and complex distributions of LLM pretraining data, along with varying training corpora for different LLMs, exacerbate the difficulty of this task. Related research is still in its early stages, with existing methods including calibration-based uncertainty estimation (Kadavath et al., 2022) and posterior approaches based on QA testing results (Yang et al., 2023), but none of them can satisfactorily solve this issue. For instances near the knowledge boundary, the situation is no longer a binary opposition between knowing and not knowing. In these cases, accurately expressing the model’s uncertainty level with fine granularity becomes a critical research question that warrants further exploration.
13.4.4. Minimizing the Alignment Tax During Hallucination Mitigation
Several studies have confirmed that the performance of ChatGPT and GPT-4 degrades over time due to continuous safety fine-tuning. This phenomenon is a typical example of an alignment tax. Similarly, when addressing hallucination problems in low-resource languages, the rate of hallucination in the original languages may increase. In the process of enhancing the faithfulness of LLMs’ multi-modal perception through visual instruction tuning on VQA datasets, the “politeness” of LLM responses may decrease due to the overly succinct nature of VQA answer annotations (Chen et al., 2023d). Furthermore, the alignment tax problem is not confined to fine-tuning-based methods but also extends to prompting-based methods such as Retrieval-Augmented Generation (RAG). These methods may also compromise the quality of responses when the quality of retrieved evidence is sub-optimal. In light of these findings, it is clear that we must explore strategies to minimize the cost during hallucination mitigation and prevent LLMs from becoming overly conservative, losing their creativity, or suffering from catastrophic forgetting in task performance.
13.4.5. Understanding Hallucination in LLMs
Although the research community has put extensive efforts to develop empirical methods for measuring and mitigating hallucinations, our understanding of LLM hallucination still remains limited. The absence of reliable theoretical frameworks and rigorous mathematical formulations hinders our ability to answer fundamental questions and make further advancements. For instance, how does hallucination correlate with model and data size (e.g., scaling laws)? What is the relationship between hallucination and auto-regressive generation? Does SFT and RLHF employ different mechanism to mitigate hallucinations? What is the lower-bound of hallucination across different tasks and domains? These questions are also intrinsically linked to a broader question – although now there are a lot of effort on empirically investigating what LLM can and cannot do from different dimensions, we still don’t understand why and how it works. There has been some attempts from various perspectives, such as knowledge encoding (Geva et al., 2021), compression (Delétang et al., 2023), model selection (Yadlowsky et al., 2023), compositionality (Dziri et al., 2023), computational perspective (Feng et al., 2023), but we are still a long way from truly understanding the mechanism behind various LLM behaviours.
14. Conclusion
In this survey, we provide the first comprehensive overview of the hallucination problem in NLG, summarizing existing evaluation metrics, mitigation methods, and the remaining challenges for future research. Hallucination is an artifact of neural-based NLG and is of concern because they appear fluent and can therefore be misleading to users. In some scenarios and tasks, hallucination can cause harm. We survey various contributors to hallucination, ranging from noisy data, erroneous parametric knowledge, incorrect attention mechanism, inappropriate training strategy, to inference exposure bias, etc. We show that there are two categories of hallucinations, namely intrinsic hallucination and extrinsic hallucination, and they need to be treated differently with diverse mitigation strategies. Hallucination is relatively easy to detect in abstractive summarization and in NMT against the evidence in the source. For dialogue systems, it is important to balance diversity vs consistency in dialogue responses. Hallucination in GQA and VL tasks is detrimental to the performance, but research on mitigation methods is still very preliminary in these areas. For data-to-text generation, hallucination arises from the discrepancy between the input and output format. Most methods to mitigate hallucinations in NMT either aim to reduce dataset noise or alleviate exposure bias. In the VL domain, models also generate unfaithful output given the visual scene, and recent works have mainly focused on the object hallucination problem. There remain many challenges ahead in identifying and mitigating hallucinations in NLG, and we hope research in this area can benefit from this survey.